### v1

# Poison-splat: Computation Cost Attack on 3D Gaussian Splatting

Abstract
3D Gaussian splatting (3DGS), known for its groundbreaking performance and efficiency, has become a dominant 3D representation and brought progress to many 3D vision tasks. However, in this work, we reveal a significant security vulnerability that has been largely overlooked in 3DGS: the computation cost of training 3DGS could be maliciously tampered by poisoning the input data. By developing an attack named Poison-splat, we reveal a novel attack surface where the adversary can poison the input images to drastically increase the computation memory and time needed for 3DGS training, pushing the algorithm towards its worst computation complexity. In extreme cases, the attack can even consume all allocable memory, leading to a Denial-of-Service (DoS) that disrupts servers, resulting in practical damages to real-world 3DGS service vendors. Such a computation cost attack is achieved by addressing a bi-level optimization problem through three tailored strategies: attack objective approximation, proxy model rendering, and optional constrained optimization. These strategies not only ensure the effectiveness of our attack but also make it difficult to defend with simple defensive measures. We hope the revelation of this novel attack surface can spark attention to this crucial yet overlooked vulnerability of 3DGS systems [^1] Our code is available at https://github.com/jiahaolu97/poison-splat..
1 Introduction
3D reconstruction, aiming to build 3D representations from multi-view 2D images, holds a critical position in computer vision and machine learning, extending its benefits across broad domains including entertainment, healthcare, archaeology, and the manufacturing industry. A notable example is Google Map with a novel feature for NeRF-based (Mildenhall et al., 2021) 3D outdoor scene reconstruction (Google, 2023). Similarly, commercial companies like  LumaAI ,  KIRI ,  Polycam and  Spline offer paid service for users to generate 3D captures from user-uploaded images or videos. Among these popular 3D applications, many are powered by a novel and sought-after 3D reconstruction paradigm known as 3D Gaussian Splatting (3DGS).
3DGS, introduced by Kerbl et al. (2023) in 2023, has quickly revolutionized the field of 3D Vision and achieved overwhelming popularity (Chen & Wang, 2024; Fei et al., 2024; Wu et al., 2024b; Dalal et al., 2024). 3DGS is not powered by neural networks, distinguishing itself from NeRF (Mildenhall et al., 2021), the former dominant force in 3D vision. Instead, it captures 3D scenes by learning a cloud of 3D Gaussians as an explicit 3D model and uses rasterization to render multiple objects simultaneously. This enables 3DGS to achieve significant advantages in rendering speed, photo-realism, and interpretability, making it a game changer in the field.
Figure 1: Our Poison-splat attack adds perturbations to input images, making 3D Gaussian Splatting need significantly more parameters to reconstruct the 3D scene, leading to huge increases in GPU memory consumption, training time and rendering latency. Here, we visualize the input image and the underlying 3D Gaussians of a clean view (left) and its corresponding poisoned view (right).
One intriguing property of Gaussian Splatting is its flexibility in model complexity. Unlike NeRF or any other neural-network based algorithms which have a pre-determined and fixed computational complexity based on network hyper-parameters, 3DGS can dynamically adjust its complexity according to the input data. During the 3DGS training process, the size of learnable parameters, i.e.the number of 3D Gaussians is dynamically adjusted to align with the complexity of the underlying 3D scene. More specifically, 3DGS algorithm uses an adaptive density control strategy to flexibly increase or decrease the number of Gaussians for optimizing reconstruction, leading to variable computational complexity in terms of GPU memory occupancy and training time cost.
The above flexibility in design aims to provide training advantages. However, this flexibility can also become a vulnerability. In this paper, we reveal a severe yet unnoticed attack vector: the flexibility of 3DGS complexity can be abused to over-consume computation resources such as GPU memory and significantly drag down training speed of Gaussian Splatting system, pushing the training process to its worst computational complexity.
We develop a computation cost attack method named Poison-splat as a proof-of-concept of this novel attack vector. Poison-splat takes the form of training data poisoning (Tian et al., 2022), where the attacker manipulates the input data fed into the victim 3DGS system. This is practical in real-world scenarios, as commercial 3D service providers like Kiri (KIRI, ), Polycam (Polycam, ) and Spline (Spline, ) receive images or videos from end users to generate 3D captures. An attacker can hide themselves among normal users to submit poisoned data and launch the attack stealthily, or even secretly manipulate data uploaded by other users. During peak usage periods, this attack competes for computing resources with legitimate users, slowing down service response time and potentially causing severe consequences like service crashes, leading to financial losses for service providers.
We formulate Poison-splat attack as a max-min problem. The inner optimization is the learning process of 3D Gaussian Splatting, i.e., minimizing the reconstruction loss given a set of input images and camera poses, while the outer optimization problem is to maximize the computation cost of solving the inner problem. Although accurately solving this bi-level optimization is often infeasible, we find that the attacker can use a proxy model to approximate the inner minimization and focus on optimizing the outer maximization objective. Moreover, we observe an important relationship that memory consumption and rendering latency exhibit a clear positive correlation with the number of 3D Gaussians in training. Therefore, the attacker can use the number of Gaussians in proxy model training as a metric to depict computation costs for outer optimization. Motivated by these insights, our Poison-splat attack uses image total variation loss as a prior to guide the over-densification of 3D Gaussians, and can approximately solve this bi-level optimization problem in a cheap way.
The contributions of this work can be concluded in three-folds:
  • We reveal that the flexibility in model complexity of 3DGS can become a security backdoor, making it vulnerable to computation cost attack. This vulnerability has been largely overlooked by the 3D vision and machine learning communities. Our research shows that such attacks are feasible, potentially causing severe financial losses to 3D service providers.
  • We formulate the attack on 3D Gaussian Splatting as a data poisoning attack problem. To our best knowledge, there is no previous work investigating how to poison training data to increase the computation cost of machine learning systems.
  • We propose a novel attack algorithm, named Poison-splat, which significantly increases GPU memory consumption and slows down training procedure of 3DGS. We hope the community can recognize this vulnerability and develop more robust 3D Gaussian Splatting algorithms or defensive methods to mitigate such attacks.
2 Preliminaries
In this section, we provide essential backgrounds on 3D Gaussian Splatting and resource-targeting attacks, which are crucial for understanding our attack method proposed in Section 3. Due to page constraints, we put related works in Appendix Section A.
2.1 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has revolutionized the 3D vision community with its superior performance and efficiency. It models a 3D scene by learning a set of 3D Gaussians $\mathcal{G}=\{G_{i}\}$ from multi-view images $\mathcal{D}=\{V_{k},P_{k}\}_{k=1}^{N}$, where each view consists of a ground-truth image ${V_{k}}$ and the corresponding camera poses ${P_{k}}$. We denote $N$ as the total number of image-pose pairs in the training data. As for the model, each 3D Gaussian $G_{i}=\{\mu_{i},s_{i},\alpha_{i},q_{i},c_{i}\}$ has 5 optimizable parameters: center coordinates $\mu_{i}$, scales $s_{i}$, opacity $\alpha_{i}$, rotation quaternion $q_{i}$, and associated view-dependent color represented as spherical harmonics $c_{i}$.
To render $\mathcal{G}$ from 3D Gaussian space to 2D images, each Gaussian is first projected into the camera coordinate frame given a camera pose $P_{k}$ to determine the depth of each Gaussian, which is the distance of that Gaussian to the camera view. The calculated depth is used for sorting Gaussians from near to far, and then the colors in 2D image are rendered in parallel by alpha blending according to the depth order of adjacent 3D Gaussians. We denote the 3DGS renderer as a function $\mathcal{R}$, which takes as input the Gaussians $\mathcal{G}$ and a camera pose $P_{k}$ to render an image $V^{\prime}_{k}=\mathcal{R}(\mathcal{G},P_{k})$.
The training is done by optimizing the reconstruction loss between the rendered image and ground truth image for each camera view, where the reconstruction loss is a combination of $\mathcal{L}_{1}$ loss and structural similarity index measure (SSIM) loss $\mathcal{L}_{\text{D-SSIM}}$ (Brunet et al., 2011):
$\mathop{\min}\limits_{\mathcal{G}}\mathcal{L}(\mathcal{D})=\mathop{\min}\limits_{\mathcal{G}}\sum\limits_{k=1}^{N}\mathcal{L}(V^{\prime}_{k},{V}_{k})=\mathop{\min}\limits_{\mathcal{G}}~{}\sum\limits_{k=1}^{N}(1-\lambda)\mathcal{L}_{1}(V^{\prime}_{k},{V}_{k})+\lambda\mathcal{L}_{\text{D-SSIM}}(V^{\prime}_{k},{V}_{k}),$ (1)
where $\lambda$ is a trade-off parameter. To ensure finer reconstruction, 3DGS deploys adaptive density control to automatically add or remove Gaussians from optimization. Adding Gaussians is called densification, while removing Gaussians is called pruning. Densification is performed when the view-space positional gradient $\nabla_{\mu_{i}}\mathcal{L}$ is larger than a pre-set gradient threshold $\tau_{g}$ (by default 0.0002). Pruning occurs when the opacity $\alpha_{i}$ falls below the opacity threshold $\tau_{\alpha}$. Both densification and pruning are non-differentiable operations. A more formal and comprehensive introduction is provided in Appendix Section B.
2.2 Resource-targeting Attack
A similar concept in computer security domain is the Denial-of-Service (DoS) attack (Elleithy et al., 2005; Aldhyani & Alkahtani, 2023; Bhatia et al., 2018). A Dos attack aims to overwhelmly consume a system’s resources or network (Mirkovic & Reiher, 2004; Long & Thomas, 2001), making it unable to serve legitimate users. Common methods include flooding the system with excessive requests or triggering a crash through malicious inputs. Such attacks pose severe risks to real-world service providers, potentially resulting in extensive operational disruptions and financial losses. For example, Midjourney, a generative AI platform, experienced a significant 24-hour system outage (Akash, 2024), potentially caused by employees from another generative AI company who were allegedly scraping data, causing a Denial-of-Service.
In the machine learning community, similar concepts are rarely mentioned. This might stem from the fact that once hyper-parameters are set in most machine learning models, such as deep neural networks, their computation complexity remains fixed. Regardless of input data content, most machine learning algorithms incur nearly constant computational costs and consume almost consistent computational resources. Despite this, only a few studies have focused on resource-targeting attacks at inference stage of machine learning systems. For example,  Shumailov et al. (2021) first identified samples that trigger excessive neuron activations which maximise energy consumption and latency. Following works investigated other inference-stage attacks against dynamic neural networks (Hong et al., 2020; Chen et al., 2023) and language models (Chen et al., 2022b; a; Gao et al., 2023). However, to our knowledge, no previous works have targeted at attacking the training-stage computation cost of machine learning systems. Our work is pioneering in bridging this research gap through Gaussian Splatting, which features adaptively flexible computation complexity.
3 Poison-splat Attack
In this section, we first formulate the computation cost attack on 3D Gaussian Splatting in Section 3.1. Subsequently, we introduce a novel attack method in Section 3.2.
3.1 Problem Formulation
We formulate our attack under data poisoning framework as follows. The victims are the service providers of 3D Gaussian Splatting (3DGS), who typically train their 3DGS models using a dataset of multi-view images and camera poses $\mathcal{D}=\{V_{k},P_{k}\}_{k=1}^{N}$. We assume the attackers have the capability to introduce poisoned data within the entire dataset. This assumption is realistic as an attacker could masquerade as a regular user, submitting poisoned data covertly, or even covertly alter data submitted by legitimate users. This allows the attack to be launched stealthily and enhances its potential impact on the training process of 3DGS. We next detail the formulation of the attacker and victim.
Attacker. The attacker begins with the clean data $\mathcal{D}=\{V_{k},P_{k}\}_{k=1}^{N}$ as input, manipulating it to produce poisoned training data $\mathcal{D}_{p}=\{\tilde{V}_{k},P_{k}\}_{k=1}^{N}$, where the attacker does not modify the camera pose configuration files. Each poisoned image $\tilde{V_{k}}$ is perturbed from the original clean image ${V_{k}}$, aiming to maximize the computational costs of victim training, such as GPU memory usage and training time. The ultimate goal of the attacker is to significantly increase the computational resources, potentially leading to a denial-of-service attack by overwhelming the training system.
Victim. On the other hand, the victim receives this poisoned dataset $\mathcal{D}_{p}=\{\tilde{V}_{k},P_{k}\}_{k=1}^{N}$ from the attacker, unaware that it has been poisoned. With this data, the victim seeks to train a Gaussian Splatting Model $\mathcal{G}$, striving to minimize the reconstruction loss (cf. Eq. (1)). The victim’s objective is to achieve the lowest loss as possible, thus ensuring the quality of the Gaussian Splatting model.
Optimization problem. To summarize, the computation cost attack for the attacker can be formulated as the following max-min bi-level optimization problem:
$\mathcal{D}_{p}=\mathop{\arg\max}\limits_{\mathcal{D}_{p}}\mathcal{C}(\mathcal{G}^{*}),~{}~{}\text{s.t.\quad}\mathcal{G}^{*}=\mathop{\arg\min}\limits_{\mathcal{G}}\mathcal{L}(\mathcal{D}_{p}),$ (2)
where the computation cost metric $\mathcal{C}(\mathcal{G})$ is flexible and can be designed by the attacker.
3.2 Proposed Method
To conduct the attack, directly addressing the above optimization is infeasible as the computation cost is not differentiable. To address this, we seek to find an approximation for the objective.
Figure 2: The figure illustrates the strong positive correlation between the number of Gaussians and (a) GPU memory occupancy, and (b) training time costs. Panel (c) shows the relationship between image sharpness, measured by the average total variation score, and the number of Gaussians.
Using the number of Gaussians as an approximation. As is mentioned in Section 1, a major advantage of 3DGS is its flexibility to dynamically adjust model complexity (i.e., the number of Gaussians) according to the complexity of the input data. This adaptability enhances the model’s efficiency and fidelity in rendering complex scenes. However, this feature may also act as a double-edged sword, providing a potential backdoor for attacks. To explore this, we analyze how the number of Gaussians affects computational costs, including memory consumption and training time. Our findings, depicted in Figures 2(a-b), reveal a clear positive correlation between computational costs and the number of Gaussians used in rendering. In light of this insight, it is intuitive to use the number of Gaussians $\|\mathcal{G}\|$ to approximate the computation cost function involved in the inner optimization:
$\mathcal{C}(\mathcal{G})\vcentcolon=\|\mathcal{G}\|.$ (3)
Maximizing the number of Gaussians by sharpening 3D objects. Even with the above approximation, solving the optimization problem is still hard since the Gaussian densification operation in 3DGS is not differentiable. Hence, it would be infeasible for the attacker to use gradient-based methods to optimize the number of Gaussians. To circumvent this, we explore a strategic alternative. As shown in Figure 2(c), we find that 3DGS tends to assign more Gaussians to those objects with more complex structures and non-smooth textures, as quantified by the total variation score—a metric assessing image sharpness (Rudin et al., 1992). Intuitively, the less smooth the surface of 3D objects is, the more Gaussians the model needs to recover all the details from its 2D image projections. Hence, non-smoothness can be a good descriptor of complexity of Gaussians, i.e., $\|\mathcal{G}\|\propto\mathcal{S}_{\text{TV}}(\mathcal{D})$. Inspired by this, we propose to maximize the computation cost by optimizing the total variation score of rendered images $\mathcal{S}_{\text{TV}}(\tilde{V}_{k})$:
$\mathcal{C}(\mathcal{G})\vcentcolon=\mathcal{S}_{\text{TV}}(\tilde{V}_{k})=\sum\limits_{i,j}\sqrt{\left|\tilde{V}_{k}^{i+1,j}-\tilde{V}_{k}^{i,j}\right|^{2}+\left|\tilde{V}_{k}^{i,j+1}-\tilde{V}_{k}^{i,j}\right|^{2}}.$ (4)
Balancing attack strength and stealthiness via optional constrained optimization. The above strategies enable the attack to significantly increase computational costs . However, this may induce an unlimited alteration to the image and result in a loss of semantic integrity in the generated view images (cf. Figure 4(b)), making it obvious to detect. Considering the strategic objectives of an adversary who may wish to preserve image semantics and launch the attack by stealth, we introduce an optional constrained optimization strategy. Inspired by adversarial attacks (Szegedy et al., 2013; Madry et al., 2018), we enforce a $\epsilon$-ball constraint of the $L_{\infty}$ norm on the perturbations:
$\tilde{V_{k}}\in\mathcal{P}_{\epsilon}(\tilde{V_{k}},V_{k}).$ (5)
where $\mathcal{P}_{\epsilon}(\cdot)$ means to clamp the rendered poisoned image to be within the $\epsilon$-ball of the $L_{\infty}$ norm around the original clean image, i.e., $\|\tilde{V_{k}}-V_{k}\|_{\infty}\leq\epsilon$. This constraint limits each pixel’s perturbation to a maximum of $\epsilon$. By tuning $\epsilon$, an attacker can balance the destructiveness and the stealthiness of the attack, allowing for strategic adjustments for the desired outcome. If $\epsilon$ is set to $\infty$, the constraint is effectively removed, returning the optimization to its original, unconstrained form.
Algorithm 1 Poison-splat
1:
Clean dataset: $\mathcal{D}=\{V_{k},P_{k}\}$; Perturbation range: $\epsilon$; Perturbation step size: $\eta$;
The iteration number of inner optimization: $T$; The iteration number of outer optimization: $\tilde{T}$.
2:Poisoned dataset: $\mathcal{D}_{p}=\{\tilde{V}_{k},P_{k}\}$
3:Initialize poisoned dataset: $\mathcal{D}_{p}=\mathcal{D}$
4:Train a Gaussian Splatting proxy model $\mathcal{G}_{p}$ on clean dataset $\mathcal{D}$triangleright Eq. (1)
5:For $t=1,..,T$ do
6: Sample a random camera view $k$
7:$V^{\prime}_{k}=\mathcal{R}(\mathcal{G}_{p},P_{k})$triangleright Get the rendered image under the camera view $k$
8:$\tilde{V}_{k}=V^{\prime}_{k}$triangleright Optimization of target image starts from rendered image
9:For $\tilde{t}=1,..,\tilde{T}$ do
10:  $\tilde{V}_{k}=\tilde{V}_{k}+\eta\cdot sign(\nabla_{\tilde{V}_{k}}\mathcal{S}_{TV}(\tilde{V}_{k}))$triangleright Eq. (4)
11:  $\tilde{V}_{k}=\mathcal{P}_{\epsilon}(\tilde{V_{k}},V_{k})$triangleright Project into $\epsilon$-ball, i.e., Eq. (5)
12:end While
13: Update proxy model $\mathcal{G}_{p}$ by minimizing $\mathcal{L}(V^{\prime}_{k},\tilde{V}_{k})$ under current view $\triangleright$ Eq. (1)
14: Update poisoned dataset: $\mathcal{D}_{p}^{k}\vcentcolon=(\tilde{V}_{k},P_{k})$triangleright Update the $k$-th element of $\mathcal{D}_{p}$
15:end While
Ensuring multi-view image consistency via proxy model. An interesting insight from our study is that simply optimizing perturbations for each view image independently by maximizing total variation score does not effectively bolster attacks. As demonstrated in Figure 3(b), this image-level TV-maximization attack is significantly less effective compared to our Poison-splat strategy. This mainly stems from the image-level optimization creating inconsistencies across the different views of poisoned images, undermining the attack’s overall effectiveness.
Our solution is inspired by the view-consistent properties of the 3DGS model’s rendering function, which effectively maintains consistency across multi-view images generated from 3D Gaussian space. In light of this, we propose to train a proxy 3DGS model to generate poisoned data. In each iteration, the attacker projects the current proxy model onto a camera pose, obtaining a rendered image $V^{\prime}_{k}$. This image then serves as a starting point for optimization, where the attacker searches for a target $\tilde{V}_{k}$ that maximizes the total variation score within the $\epsilon$-bound of the clean image $V_{k}$. Following this, the attacker updates the proxy model with a single optimization step to mimic the victim’s behavior. In subsequent iterations, the generation of poisoned images initiates with the rendering output from the updated proxy model. In this way, the attacker approximately solves this bi-level optimization by iteratively unrolling the outer and inner optimization, enhancing the attack’s effectiveness while maintaining consistency across views. We summarize the pipeline of Poison-splat in Algorithm 1.
4 Experiments
Datasets. All experiments in this paper are carried out on three common 3D datasets: (1) NeRF-Synthetic [^2] Dataset publicly accessible at https://github.com/bmild/nerf. (Mildenhall et al., 2021) is a synthetic 3D object dataset, which contains multi-view renderings of synthetic 3D objects. (2) Mip-NeRF360333Dataset publicly accessible at https://jonbarron.info/mipnerf360/ dataset (Barron et al., 2022) is a 3D scene dataset, which is composed of photographs of large real-world outdoor scenes. (3) Tanks-and-Temples [^4] Dataset publicly accessible at https://www.tanksandtemples.org/download/ dataset (Knapitsch et al., 2017) contains realistic 3D captures including both outdoor scenes and indoor environments.
Implementation details. In our experiments, we use the official implementation of 3D Gaussian Splatting [^5] https://github.com/graphdeco-inria/gaussian-splatting (Kerbl et al., 2023) to implement the victim behavior. Following their original implementation, we use the recommended default hyper-parameters, which were proved effective across a broad spectrum of scenes from various datasets. For testing black-box attack effectiveness, we use the official implementation of Scaffold-GS [^6] https://github.com/city-super/Scaffold-GS (Lu et al., 2024) as a black box victim. All experiments reported in the paper were carried out on a single NVIDIA A800-SXM4-80G GPU.
Evaluation metrics. We use the number of 3D Gaussians, GPU memory occupancy and the training time cost as metrics to evaluate the computational cost of 3DGS. A more successful attack is characterized by larger increases in the number of Gaussians, higher GPU memory consumption, and extended training time, compared with training on clean data. We also report the rendering speed of the resulted 3DGS models to show the increase in model complexity and degradation of inference latency as a byproduct of our attack.
(a) Overall comparisons of computation cost in training time and GPU memory occupancy on MIP-NeRF360 (room).
(b) Number of Gaussians needed for reconstruction under our attack, naive TV-maximization attack and on the clean data of MIP-NeRF360 (room).
Figure 3: Figure (a) shows that our attack leads to a significant increase on the GPU memory and training time. Figure (b) shows that attacks simply maximizing total variation score at the image level are less effective compared to Poison-splat, which highlights the crucial role of the proxy model in our attack design.
4.1 Main Results
We report the comparisons of the number of Gaussians, peak GPU memory, training time and rendering speed between clean and poisoned data across three different datasets in Tables 1 to show the effectiveness of our attack. The full results containing all scenes are listed in the Appendix Table 3 and Table 4. We provide visualizations of example poisoned datasets and corresponding victim reconstructions in Appendix Section E.
We want to highlight some observations from the results. First, our Poison-splat attack demonstrates the ability to craft a huge extra computational burden across multiple datasets. Even with perturbations constrained within a small range in an $\epsilon$-constrained attack, the peak GPU memory can be increased to over 2 times, making the overall maximum GPU occupancy higher than 24 GB - in the real world, this may mean that our attack may require more allocable resources than common GPU stations can provide, e.g., RTX 3090, RTX 4090 and A5000. Furthermore, as we visualized the computational cost in Figure 3(a), the attack not only significantly increases the memory usage, but also greatly slows down training speed. This property would further strengthen the attack, since the overwhelming GPU occupancy will last longer than normal training may take, making the overall loss of computation power higher. As shown in the bottom of Table 1, if we do not put constraints on the attacker and let it modify the image pixels unlimitedly, then the attack results would be even more damaging. For instance, in the NeRF-Synthetic (ship) scene, the peak GPU usage could increase up to 21 times and almost reach 80 GB GPU memory. This level approaches the upper limits of the most advanced GPUs currently available, such as H100 and A100, and is undeniably destructive.
Table 1: Attack results of Poison-splat across multiple datasets. Constrained attacks ($\epsilon=16/255$) can covertly increase training costs through minor perturbations to images. In contrast, unconstrained attacks allow for unlimited modifications to the original input, leading to significantly higher resource consumption and more extensive damage. We highlight attacks which successfully consume over 24 GB GPU memory, which can cause a denial-of-service(i.e., out-of-memory error and service halting) if 3DGS service is provided on a common 24GB-memory GPU (e.g., RTX 3090, RTX 4090 or A5000). Full results on all scenes provided in Table 3 and Table 4.
Constrained Poison-splat attack with $\epsilon=16/255$
Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes] Render speed [FPS]
Scene Setting clean poisoned clean poisoned clean poisoned clean poisoned
NS-hotdog 0.185 M 1.147 M (6.20$\uparrow$) 3336 29747 (8.92$\uparrow$) 8.57 17.43 (2.03$\uparrow$) 443 69 (6.42$\downarrow$)
NS-lego 0.341 M 0.805 M (2.36$\uparrow$) 3532 9960 (2.82$\uparrow$) 8.62 13.02 (1.51$\uparrow$) 349 145 (2.41$\downarrow$)
NS-materials 0.169 M 0.410 M (2.43$\uparrow$) 3172 4886 (1.54$\uparrow$) 7.33 9.92 (1.35$\uparrow$) 447 250 (1.79$\downarrow$)
NS-ship 0.272 M 1.071 M (3.94$\uparrow$) 3692 16666 (4.51$\uparrow$) 8.87 16.12 (1.82$\uparrow$) 326 87 (3.75$\downarrow$)
TT-Courthouse 0.604 M 0.733 M (1.21$\uparrow$) 11402 12688 (1.11$\uparrow$) 13.81 14.49 (1.05$\uparrow$) 284 220 (1.29$\downarrow$)
TT-Courtroom 2.890 M 5.450 M (1.89$\uparrow$) 9896 16308 (1.65$\uparrow$) 16.88 23.92 (1.42$\uparrow$) 139 83 (1.67$\downarrow$)
TT-Museum 4.439 M 7.130 M (1.61$\uparrow$) 12704 18790 (1.48$\uparrow$) 19.79 26.38 (1.33$\uparrow$) 121 78 (1.55$\downarrow$)
TT-Playground 2.309 M 4.307 M (1.87$\uparrow$) 8717 13032 (1.50$\uparrow$) 15.17 21.82 (1.44$\uparrow$) 151 87 (1.74$\downarrow$)
MIP-bicycle 5.793 M 10.129 M (1.75$\uparrow$) 17748 27074 (1.53$\uparrow$) 33.42 44.44 (1.33$\uparrow$) 69 43 (1.60$\downarrow$)
MIP-counter 1.195 M 4.628 M (3.87$\uparrow$) 10750 29796 (2.77$\uparrow$) 26.46 37.63 (1.42$\uparrow$) 164 51 (3.22$\downarrow$)
MIP-room 1.513 M 7.129 M (4.71$\uparrow$) 12316 46238 (3.75$\uparrow$) 25.36 47.98 (1.89$\uparrow$) 154 35 (4.40$\downarrow$)
MIP-stump 4.671 M 10.003 M (2.14$\uparrow$) 14135 25714 (1.82$\uparrow$) 27.06 38.48 (1.42$\uparrow$) 111 57 (1.95$\downarrow$)
Unconstrained Poison-splat attack with $\epsilon=\infty$
Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes] Render speed [FPS]
Scene Setting clean poisoned clean poisoned clean poisoned clean poisoned
NS-hotdog 0.185 M 4.272 M (23.09$\uparrow$) 3336 47859 (14.35$\uparrow$) 8.57 38.85 (4.53$\uparrow$) 443 29 (15.28$\downarrow$)
NS-lego 0.341 M 4.159 M (12.20$\uparrow$) 3532 78852 (22.33$\uparrow$) 8.62 42.46 (4.93$\uparrow$) 349 25 (13.96$\downarrow$)
NS-mic 0.205 M 3.940 M (19.22$\uparrow$) 3499 61835 (17.67$\uparrow$) 8.08 39.02 (4.83$\uparrow$) 300 29 (10.34$\downarrow$)
NS-ship 0.272 M 4.317 M (15.87$\uparrow$) 3692 80956 (21.93$\uparrow$) 8.87 44.11 (4.97$\uparrow$) 326 24 (13.58$\downarrow$)
TT-Courthouse 0.604 M 3.388 M (5.61$\uparrow$) 11402 29856 (2.62$\uparrow$) 13.81 25.33 (1.83$\uparrow$) 284 54 (5.26$\downarrow$)
TT-Courtroom 2.890 M 13.196 M (4.57$\uparrow$) 9896 33871 (3.42$\uparrow$) 16.88 41.69 (2.47$\uparrow$) 139 41 (3.39$\downarrow$)
TT-Museum 4.439 M 16.501 M (3.72$\uparrow$) 12704 43317 (3.41$\uparrow$) 19.79 48.89 (2.47$\uparrow$) 121 36 (3.36$\downarrow$)
TT-Playground 2.309 M 10.306 M (4.46$\uparrow$) 8717 27304 (3.13$\uparrow$) 15.17 38.77 (2.56$\uparrow$) 151 39 (3.87$\downarrow$)
MIP-bicycle 5.793 M 25.268 M (4.36$\uparrow$) 17748 63236 (3.56$\uparrow$) 33.42 81.48 (2.44$\uparrow$) 69 16 (4.31$\downarrow$)
MIP-counter 1.195 M 11.167 M (9.34$\uparrow$) 10750 80732 (7.51$\uparrow$) 26.46 62.04 (2.34$\uparrow$) 164 19 (8.63$\downarrow$)
MIP-room 1.513 M 16.019 M (10.59$\uparrow$) 12316 57540 (4.67$\uparrow$) 25.36 76.25 (3.01$\uparrow$) 154 17 (9.06$\downarrow$)
MIP-stump 4.671 M 13.550 M (2.90$\uparrow$) 14135 36181 (2.56$\uparrow$) 27.06 51.51 (1.90$\uparrow$) 111 27 (4.11$\downarrow$)
(a) Evolution of the proxy model in an $\epsilon$-constrained attack
(b) Evolution of the proxy model in an unconstrained attack
Figure 4: Visualizations of proxy model updates during an attack process. The proxy 3DGS model gradually obtains more complexity from learning from non-smoothness in the 2D image space.
To help the readers build more understanding into the attack procedure, we visualize the internal states during the attack optimization procedure, rendering the proxy model in Figure 4. It is evident that the proxy model can be guided from non-smoothness of 2D images to develop highly complex 3D shapes. Consequently, the poisoned data produced from the projection of this over-densified proxy model can produce more poisoned data, inducing more Gaussians to fit these poisoned data.
4.2 Black-box Attack Performance
The attack results described previously assume a white-box scenario, where the attacker is fully aware of the victim system’s details. In real-world, companies may utilize proprietary algorithms of their own, where the attacker may have no knowledge about their hyper-parameter settings or even their underlying 3DGS representations. This raises an important question: Can Poison-splat remain effective against black-box victim systems?
The answer is affirmative. To test the effectiveness of our attack on a black-box victim, we chose Scaffold-GS (Lu et al., 2024), a 3DGS variant algorithm designed with a focus on efficiency. Scaffold-GS differs significantly from traditional Gaussian Splatting as it utilizes anchor points to distribute local 3D Gaussians, which greatly alters the underlying feature representation. We directly tested Scaffold-GS on the poisoned dataset we crafted for the traditional 3DGS algorithm, making the victim (Scaffold-GS trainer) a black-box model. The results, as presented in Table 2 and Table 5, demonstrate the attack’s capability to generalize effectively, as evidenced by increased Gaussian parameters, higher GPU memory consumption, and reduced training speed.
The key reason for the success of generalization ability to black-box victims is our core intuition: promoting the non-smoothness of input images while maintaining their 3D consistency. This approach ensures that our poisoned dataset effectively increases the number of 3D-representation units for all adaptive complexity 3D reconstruction algorithms, regardless of the specific algorithm used.
Table 2: Black-box attack results on NeRF-Synthetic and MIP-NeRF360 dataset. The victim system, Scaffold-GS, utilizes distinctly different Gaussian Splatting feature representations compared to traditional Gaussian Splatting and remains unknown to the attacker. These results demonstrate the robust generalization ability of our attack against unknown black-box victim systems. Full results provided in Table 5.
Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes]
Scene Setting clean poisoned clean poisoned clean poisoned
Constrained Black-box Poison-splat attack with $\epsilon=16/255$ against Scaffold-GS
NS-lego 0.414 M 2.074 M (5.01$\uparrow$) 3003 4808 (1.60$\uparrow$) 9.77 17.91 (1.83$\uparrow$)
NS-ship 1.000 M 3.291 M (3.29$\uparrow$) 3492 5024 (1.44$\uparrow$) 11.68 21.92 (1.88$\uparrow$)
MIP-bonsai 4.368 M 10.608 M (2.43$\uparrow$) 10080 12218 (1.21$\uparrow$) 35.33 37.71 (1.07$\uparrow$)
MIP-stump 6.798 M 14.544 M (2.14$\uparrow$) 7322 9432 (1.29$\uparrow$) 33.53 36.32 (1.08$\uparrow$)
Unconstrained Black-box Poison-splat attack with $\epsilon=\infty$ against Scaffold-GS
NS-lego 0.414 M 3.973 M (9.60$\uparrow$) 3003 6242 (2.08$\uparrow$) 9.77 26.11 (2.67$\uparrow$)
NS-ship 1.000 M 4.717 M (4.72$\uparrow$) 3492 6802 (1.95$\uparrow$) 11.68 28.22 (2.42$\uparrow$)
MIP-bonsai 4.368 M 28.042 M (6.42$\uparrow$) 10080 22115 (2.19$\uparrow$) 35.33 78.36 (2.22$\uparrow$)
MIP-stump 6.798 M 34.027 M (5.01$\uparrow$) 7322 20797 (2.84$\uparrow$) 33.53 79.64 (2.38$\uparrow$)
4.3 Ablation Study on proxy model
We next ablate the effectiveness of our proxy model. As shown in Figure 3(b), we compare the number of Gaussians produced in victim learning procedures with two baselines: the green curve represents trajectory training on clean data, and the yellow curve represents a naive Total Variation maximizing attack, which does not involve a proxy model, but only maximize Total Variation score on each view individually. As we show in Figure 3(b), this naive attack is not as effective as our attack design. We attribute this reason to be lack of consistency of multi-view information. Since the Total Variation loss is a local image operator, individually maximizing the TV loss on each image will result in a noised dataset without aligning with each other. As a result, victim learning from this noised set of images struggles to align shapes in 3D space, and the training on fine details becomes harder to converge. Thus, when faced with conflicting textures from different views, the Gaussian Splatting technique fails to effectively perform reconstruction on details, leading to fewer Gaussians produced compared with our attack. We use this study to highlight the necessity of decoy model in our attack methodology design.
4.4 Discussions on Naive Defense Strategies
Our attack method maximizes the computational costs of 3DGS by optimizing the number of Gaussians. One may argue that a straightforward defense method is to simply impose an upper limit on the number of Gaussians during victim training to thwart the attack. However, such a straightforward strategy is not as effective as expected. While limiting the number of Gaussians might mitigate the direct impact on computational costs, it significantly degrades the quality of 3D reconstruction, as demonstrated in Figure 5. The figure shows that 3DGS models trained under these defensive constraints perform much worse compared to those with unconstrained training, particularly in terms of detail reconstruction. This decline in quality occurs because 3DGS cannot automatically distinguish necessary fine details from poisoned textures. Naively capping the number of Gaussians will directly lead to the failure of the model to reconstruct the 3D scene accurately, which violates the primary goal of the service provider. This study demonstrates more sophisticated defensive strategies are necessary to both protect the system and maintain the quality of 3D reconstructions under our attack.
Figure 5: Poison-splat attack cannot be painlessly defended by simply constraining the number of Gaussians during 3DGS optimization. As demonstrated, while this defense can cap the maximum resource consumption, it markedly degrades the quality of 3D reconstruction, in terms of photo-realism and fine-grained details of the rendered images, which violates the primary goal of the service provider.
5 Conclusion and Discussions
This paper uncovers a significant yet previously overlooked security vulnerability in 3D Gaussian Splatting (3DGS). To this end, we have proposed a novel Poison-splat attack method, which dramatically escalates the computational demands of 3DGS and can even trigger a Denial-of-Service (e.g., server disruption), causing substantial financial losses for 3DGS service providers. By leveraging a sophisticated bi-level optimization framework and strategies such as attack objective approximation, proxy model rendering, and optional constrained optimization, we demonstrate the feasibility of such attacks and emphasize the challenges in countering them with naive defenses. To the best of our knowledge, this is the first research work to explore an attack on 3DGS, and the first attack targeting the computation complexity of training phase in machine learning systems. We encourage researchers and practitioners in the field of 3DGS to acknowledge this security vulnerability and work together towards more robust algorithms and defense strategies against such threats. Lastly, we would like to discuss potential limitations, future directions, and social impacts of our work.
Limitations and Future directions. There are still limitations of our work, and we expect future studies can work on them:
  1. 1.
    Better approximation of outer maximum optimization. In this work, we approximate the outer maximum objective (i.e., the computational cost) with the number of Gaussians. Although the number of Gaussians strongly correlates with GPU memory occupancy and rendering latency, there could be better optimizable metrics for this purpose. For example, the "density" of Gaussians, i.e., the number of Gaussians falling into the same tile to involve in alpha-blending, might be a better metric to achieve better optimization outcomes.
  2. 2.
    Better defense methods. We focus on developing attack methods, without delving deeply into defensive strategies. We hope future studies can propose more robust 3DGS algorithms or develop more effective defensive techniques to counteract such attacks. This focus could significantly enhance the security and reliability of 3DGS systems in practical applications.
Societal Impacts. While our methods could potentially be misused by malicious actors to disrupt 3DGS service providers and cause financial harm, our aim is not to facilitate such actions. Instead, our objective is to highlight significant security vulnerabilities within 3DGS systems and prompt the community—researchers, practitioners, and providers—to recognize and address these gaps. We aim to inspire the development of more robust algorithms and defensive strategies, enhancing the security and reliability of 3DGS systems in practical applications. We are committed to ethical research and do not endorse using our findings to inflict societal harm.
6 Acknowledgement
We sincerely thank Xuanyu Yi (NTU), Xinjie Zhang (HKUST), Jiaqi Ma (WHU) and Shizun Wang (NUS) for their helpful discussions.
Appendix
The Appendix is organized as follows: Section A discusses related works; Section B provides a formal and comprehensive description of 3D Gaussian Splatting training. Section C offers the full results of experiments mentioned in the main paper, presented here due to page constraints. Section D lists all the notations used throughout the paper. Section E presents examples of attacker-crafted poisoned images and the corresponding reconstructions from victim models trained with those images.
Appendix A Related Works
A.1 3D Gaussian Splatting
3D Gaussian Splatting is receiving overwhelming popularity in the 3D vision domain (Chen & Wang, 2024; Fei et al., 2024; Wu et al., 2024b; Dalal et al., 2024). The explicit nature of 3D Gaussian Splatting enables real-time rendering capabilities and unparalleled levels of control and editability, demonstrating its effectiveness in 3D generation. Since its introduction in (Kerbl et al., 2023), 3D Gaussian Splatting has facilitated a diverse array of downstream applications, including text-to-3D generation (Chen et al., 2024; Tang et al., 2024; Yi et al., 2024), Simultaneous Localization and Mapping (Keetha et al., 2024; Matsuki et al., 2024; Yan et al., 2024), human avatars (Kocabas et al., 2024; Lei et al., 2024; Zielonka et al., 2023), 4D-dynamic scene modeling (Luiten et al., 2024; Wu et al., 2024a; Yang et al., 2024), and other applications (Xie et al., 2024b; Shen et al., 2024; Ye et al., 2023; Wang et al., 2024b; a).
While the majority of the audience in the 3D vision community places their focus on utility, we found only a handful of research works put efforts into investigating the security issues of 3D algorithms. Xie et al. (2024a) analyzed the resilience of 3D object detection systems under white-box and black-box attacks. (Fu et al., 2023; Horváth & Józsa, 2023) proposed attacks against Neural Radiance Fields (NeRFs). To our knowledge, this is the first attack ever proposed for 3D Gaussian Splatting systems.
A.2 Attacks on Machine Learning Systems
Machine learning has been shown to be vulnerable to various types of attacks (Pitropakis et al., 2019; Rigaki & Garcia, 2023; Chakraborty et al., 2018; Oliynyk et al., 2023; Tian et al., 2022). The majority of attacks target the either confidentiality (including membership inference (Shokri et al., 2017; Salem et al., 2019; Choquette-Choo et al., 2021), data reconstruction attacks (Fredrikson et al., 2015; Carlini et al., 2021; 2023; Yu et al., 2023; Geiping et al., 2020; Lu et al., 2022) and model stealing attacks (Tramèr et al., 2016; Chandrasekaran et al., 2020)) or integrity (like adversarial attacks (Biggio et al., 2013; Szegedy et al., 2013; Dong et al., 2018; 2019) and data poisoning attacks (Barreno et al., 2010; Jagielski et al., 2018; Biggio et al., 2012; Mei & Zhu, 2015b)). Different from the above mentioned attacks, we aim to illustrate that by carefully crafting input images, the attacker can indeed attack availability, i.e., timely and cost-affordable access to machine learning service. The most similar research is the work by Shumailov et al. (2021), which is the first Denial-of-Service (DoS) attack towards machine learning systems. It proposed sponge examples which can cause much more latency and energy consumption for language models and image classifiers, driving deep learning models to their worst performance. Different from the work (Shumailov et al., 2021), our attack targets 3D reconstruction algorithms and is not driven by neural networks.
A.2.1 Data Poisoning Attack
We consider our proposed attacks to fall into the category of data poisoning attack, which manipulates training data to launch the attack at the training stage. The attacker makes the first move to craft poisoned training samples, and the victim has a goal of training the system based on samples provided by the attacker. Solving this problem typically involves solving a max-min bi-level optimization problem, where the attacker wants to maximize some adversarial loss (i.e., test error or computational cost) and victim wants to minimize the training objective (i.e., training loss). Different literature solves the problem using different techniques (Biggio et al., 2012; Mei & Zhu, 2015b; a; Li et al., 2016). For many machine learning systems, given the inherent complexity and their reliance on large datasets, it is intractable to exactly do the bi-level optimization. Muñoz-González et al. (2017) exploited back-gradient optimization technique to approximate real gradients through reverse-mode differentiation.  Huang et al. (2020) formulated as a meta-learning problem, and approximated the inner optimal solution by training only $K$ SGD steps for each outer objective evaluation. Koh et al. proposed three methods to approximate the expensive bi-level optimization: influence functions (Koh & Liang, 2017), minimax duality and Karush-Kuhn-Tucker (KKT) conditions. Fowl et al. (2021a) used gradient matching which assumes inner optimality has been achieved by a fixed model. Fowl et al. (2021b) avoided the bi-level optimization with a well-trained proxy model from clean data. Different from these previous studies, we propose a gradient-free optimization process to solve the outer optimization pipeline, and show a good optimization result on the attacker’s goal.
A.2.2 Energy-Latency Attacks in Machine Learning
A limited number of studies have investigated attacks to manipulate the computation cost needed for machine learning systems, mainly in inference-stage. Shumailov et al. (2021) initially brought people’s attention to sponge samples, which are malicious samples able to maximize the energy consumption and responsive latency of a neural network. Following this work, there are other works studying samples which can cause the worst case computation resource consumption in a (dynamic) machine learning systems, for example Hong et al. (2020) and Chen et al. (2023) found samples that can necessitate late exit in a dynamic-depth neural network. More up-to-date works utilize the flexible nature of language to attack NLP models, inducing extended natural language sequence responses to over-consume energy and latency (Chen et al., 2022b; a; Gao et al., 2023). However, to the best of our knowledge, there has been no previous work targeting Denial-of-Service attacks at the training stage of machine learning systems. Our work is pioneering in proposing a DoS attack specifically designed for the training phase of machine learning systems. Additionally, this is also the first effort to develop an attack against 3D Gaussian Splatting.
Appendix B More Backgrounds of 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has revolutionized the 3D vision community with its superior performance and efficiency. It models a 3D scene by learning a set of 3D Gaussians $\mathcal{G}=\{G_{i}\}$ from multi-view images $\mathcal{D}=\{V_{k},P_{k}\}_{k=1}^{N}$, where each view consists of a ground-truth image ${V_{k}}$ and the corresponding camera poses ${P_{k}}$. We denote $N$ as the total number of image-pose pairs in the training data, which can be arbitrary and unfixed in 3DGS. As for the model, each 3D Gaussian $G_{i}=\{\mu_{i},s_{i},\alpha_{i},q_{i},c_{i}\}$ has 5 optimizable parameters: center coordinates $\mu_{i}\in\mathbb{R}^{3}$, scales $s_{i}\in\mathbb{R}^{3}$, opacity $\alpha_{i}\in\mathbb{R}$, rotation quaternion $q_{i}\in\mathbb{R}^{4}$, and associated view-dependent color represented as spherical harmonics $c_{i}\in\mathbb{R}^{3(d+1)^{2}}$, where $d$ is the degree of spherical harmonics.
To render $\mathcal{G}$ from 3D Gaussian space to 2D images, each Gaussian is first projected into the camera coordinate frame given a camera pose $P_{k}$ to determine the depth of each Gaussian, which is the distance of that Gaussian to the camera view. The calculated depth is used for sorting Gaussians from near to far, and then the colors in 2D image are rendered in parallel by alpha blending according to the depth order of adjacent 3D Gaussians. The core reason for 3DGS to achieve real-time optimization and rendering lies in this tile-based differentiable rasterization pipeline. We denote the 3DGS renderer as a function $\mathcal{R}$, which takes as input the Gaussians $\mathcal{G}$ and a camera pose $P_{k}$ to render an image $V^{\prime}_{k}=\mathcal{R}(\mathcal{G},P_{k})$.
The training is done by optimizing the reconstruction loss between the rendered image and ground truth image for each camera view, where the reconstruction loss is a combination of $\mathcal{L}_{1}$ loss and structural similarity index measure (SSIM) loss $\mathcal{L}_{\text{D-SSIM}}$:
$\mathop{\min}\limits_{\mathcal{G}}\mathcal{L}(\mathcal{D})=\mathop{\min}\limits_{\mathcal{G}}\sum\limits_{k=1}^{N}\mathcal{L}(V^{\prime}_{k},{V}_{k})=\mathop{\min}\limits_{\mathcal{G}}~{}\sum\limits_{k=1}^{N}(1-\lambda)\mathcal{L}_{1}(V^{\prime}_{k},{V}_{k})+\lambda\mathcal{L}_{\text{D-SSIM}}(V^{\prime}_{k},{V}_{k}),$ (6)
where $\lambda$ is a trade-off parameter. To ensure finer reconstruction, Gaussian Splatting deploys adaptive density control to automatically add or remove Gaussians from optimization. The process of adding Gaussians, known as densification, involves two primary operations: splitting and cloning. Conversely, removing Gaussians, called pruning, indicates eliminating a Gaussian from the set of optimization variables. Both densification and pruning are non-differentiable operations. The adaptive density control is performed during regular optimization intervals and follows specific, predefined rules. In the initial design (Kerbl et al., 2023), for a Gaussian $G_{i}$ with view-space projection point $\mu^{k}_{i}=(\mu^{k}_{i,x},\mu^{k}_{i,y})$ under viewpoint $V_{k}$, the average view-space positional gradient $\nabla_{\mu_{i}}\mathcal{L}$ is calculated every 100 training iterations:
$\nabla_{\mu_{i}}\mathcal{L}=\frac{\sum_{k=1}^{M}\|\frac{\partial\mathcal{L}(V^{\prime}_{k},{V}_{k})}{\partial\mu_{i}^{k}}\|}{M},$ (7)
where $M$ is the total number of views that the Gaussian $G_{i}$ participates in calculation during the 100 iterations (Ye et al., 2024) and $M$ is usually less than $N$. Densification is performed when the view-space positional gradient $\nabla_{\mu_{i}}\mathcal{L}$ is larger than a pre-set gradient threshold $\tau_{g}$ (by default 0.0002). Moreover, the decision to split or clone a Gaussian $G_{i}$ is based on its scale vector $s_{i}$, which is compared with a scale threshold $\tau_{s}$: it is split if larger, and cloned if smaller. Pruning occurs when the opacity $\alpha_{i}$ falls below the opacity threshold $\tau_{\alpha}$. We also notice some recent 3DGS studies (Cheng et al., 2024; Fang & Wang, 2024; Li et al., 2024; Bulò et al., 2024) that adopt different yet still pre-defined densification rules.
Appendix C Extended Experiments
C.1 Full attack results
We report the constrained attack results with $\epsilon=16/255$ on all scenes of NeRF-Synthetic, Mip-NeRF360 and Tanks-and-Temples dataset in Table 3, and the unconstrained attack results in Table 4.
Table 3: Full constrained attack results ($\epsilon=16/255$) on all scenes from the NeRF-Synthetic, MIP-NeRF360 and Tanks-and-Temples datasets. All results are reported by averaging over three individual runs.
Constrained Poison-splat attack with $\epsilon=16/255$
Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes] Render speed [FPS]
Scene Setting clean poisoned clean poisoned clean poisoned clean poisoned
NS-chair 0.494 M 0.957 M (1.94$\uparrow$) 3752 8003 (2.13$\uparrow$) 9.45 13.41 (1.42$\uparrow$) 260 117 (2.22$\downarrow$)
NS-drums 0.385 M 0.695 M (1.81$\uparrow$) 3456 7190 (2.08$\uparrow$) 8.75 11.94 (1.36$\uparrow$) 347 144 (2.41$\downarrow$)
NS-ficus 0.265 M 0.273 M (1.03$\uparrow$) 3194 3783 (1.18$\uparrow$) 6.99 8.94 (1.28$\uparrow$) 346 337 (1.03$\downarrow$)
NS-hotdog 0.185 M 1.147 M (6.20$\uparrow$) 3336 29747 (8.92$\uparrow$) 8.57 17.43 (2.03$\uparrow$) 443 69 (6.42$\downarrow$)
NS-lego 0.341 M 0.805 M (2.36$\uparrow$) 3532 9960 (2.82$\uparrow$) 8.62 13.02 (1.51$\uparrow$) 349 145 (2.41$\downarrow$)
NS-materials 0.169 M 0.410 M (2.43$\uparrow$) 3172 4886 (1.54$\uparrow$) 7.33 9.92 (1.35$\uparrow$) 447 250 (1.79$\downarrow$)
NS-mic 0.205 M 0.359 M (1.75$\uparrow$) 3499 6437 (1.84$\uparrow$) 8.08 10.96 (1.36$\uparrow$) 300 172 (1.74$\downarrow$)
NS-ship 0.272 M 1.071 M (3.94$\uparrow$) 3692 16666 (4.51$\uparrow$) 8.87 16.12 (1.82$\uparrow$) 326 87 (3.75$\downarrow$)
MIP-bicycle 5.793 M 10.129 M (1.75$\uparrow$) 17748 27074 (1.53$\uparrow$) 33.42 44.44 (1.33$\uparrow$) 69 43 (1.60$\downarrow$)
MIP-bonsai 1.294 M 6.150 M (4.75$\uparrow$) 10216 20564 (2.01$\uparrow$) 21.68 32.51 (1.50$\uparrow$) 214 66 (3.24$\downarrow$)
MIP-counter 1.195 M 4.628 M (3.87$\uparrow$) 10750 29796 (2.77$\uparrow$) 26.46 37.63 (1.42$\uparrow$) 164 51 (3.22$\downarrow$)
MIP-flowers 3.508 M 5.177 M (1.48$\uparrow$) 12520 16338 (1.30$\uparrow$) 25.47 30.05 (1.18$\uparrow$) 124 89 (1.39$\downarrow$)
MIP-garden 5.684 M 7.697 M (1.35$\uparrow$) 17561 23115 (1.32$\uparrow$) 33.19 40.82 (1.23$\uparrow$) 81 54 (1.50$\downarrow$)
MIP-kitchen 1.799 M 6.346 M (3.53$\uparrow$) 12442 20902 (1.68$\uparrow$) 26.61 43.24 (1.62$\uparrow$) 133 46 (2.89$\downarrow$)
MIP-room 1.513 M 7.129 M (4.71$\uparrow$) 12316 46238 (3.75$\uparrow$) 25.36 47.98 (1.89$\uparrow$) 154 35 (4.40$\downarrow$)
MIP-stump 4.671 M 10.003 M (2.14$\uparrow$) 14135 25714 (1.82$\uparrow$) 27.06 38.48 (1.42$\uparrow$) 111 57 (1.95$\downarrow$)
MIP-treehill 3.398 M 5.684 M (1.67$\uparrow$) 11444 17287 (1.51$\uparrow$) 22.76 29.26 (1.29$\uparrow$) 123 70 (1.76$\downarrow$)
TT-Auditorium 0.699 M 2.719 M (3.89$\uparrow$) 5982 14532 (2.43$\uparrow$) 12.69 19.05 (1.50$\uparrow$) 266 100 (2.66$\downarrow$)
TT-Ballroom 3.103 M 3.957 M (1.28$\uparrow$) 10416 12226 (1.17$\uparrow$) 20.37 22.43 (1.10$\uparrow$) 96 85 (1.13$\downarrow$)
TT-Barn 1.005 M 1.865 M (1.86$\uparrow$) 7813 10470 (1.34$\uparrow$) 13.10 16.27 (1.24$\uparrow$) 229 127 (1.80$\downarrow$)
TT-Caterpillar 1.290 M 1.857 M (1.44$\uparrow$) 6624 8014 (1.21$\uparrow$) 11.96 14.77 (1.23$\uparrow$) 228 164 (1.39$\downarrow$)
TT-Church 2.351 M 3.396 M (1.44$\uparrow$) 10076 11797 (1.17$\uparrow$) 17.43 20.54 (1.18$\uparrow$) 132 97 (1.36$\downarrow$)
TT-Courthouse 0.604 M 0.733 M (1.21$\uparrow$) 11402 12688 (1.11$\uparrow$) 13.81 14.49 (1.05$\uparrow$) 284 220 (1.29$\downarrow$)
TT-Courtroom 2.890 M 5.450 M (1.89$\uparrow$) 9896 16308 (1.65$\uparrow$) 16.88 23.92 (1.42$\uparrow$) 139 83 (1.67$\downarrow$)
TT-Family 2.145 M 3.734 M (1.74$\uparrow$) 7261 10707 (1.47$\uparrow$) 13.58 18.46 (1.36$\uparrow$) 182 111 (1.64$\downarrow$)
TT-Francis 0.759 M 1.612 M (2.12$\uparrow$) 5208 6938 (1.33$\uparrow$) 10.19 13.92 (1.37$\uparrow$) 279 168 (1.66$\downarrow$)
TT-Horse 1.308 M 2.510 M (1.92$\uparrow$) 5125 7762 (1.51$\uparrow$) 12.05 16.47 (1.37$\uparrow$) 220 121 (1.82$\downarrow$)
TT-Ignatius 3.144 M 3.937 M (1.25$\uparrow$) 9943 11800 (1.19$\uparrow$) 16.32 18.52 (1.13$\uparrow$) 149 118 (1.26$\downarrow$)
TT-Lighthouse 0.888 M 1.178 M (1.33$\uparrow$) 7466 8696 (1.16$\uparrow$) 14.21 15.92 (1.12$\uparrow$) 213 163 (1.31$\downarrow$)
TT-M60 1.667 M 2.980 M (1.79$\uparrow$) 7948 10874 (1.37$\uparrow$) 14.08 18.22 (1.29$\uparrow$) 200 107 (1.87$\downarrow$)
TT-Meetingroom 1.267 M 2.927 M (2.31$\uparrow$) 6645 10234 (1.54$\uparrow$) 13.53 18.97 (1.40$\uparrow$) 197 102 (1.93$\downarrow$)
TT-Museum 4.439 M 7.130 M (1.61$\uparrow$) 12704 18790 (1.48$\uparrow$) 19.79 26.38 (1.33$\uparrow$) 121 78 (1.55$\downarrow$)
TT-Panther 1.812 M 3.457 M (1.91$\uparrow$) 8511 12048 (1.42$\uparrow$) 13.49 19.46 (1.44$\uparrow$) 199 103 (1.93$\downarrow$)
TT-Playground 2.309 M 4.307 M (1.87$\uparrow$) 8717 13032 (1.50$\uparrow$) 15.17 21.82 (1.44$\uparrow$) 151 87 (1.74$\downarrow$)
TT-Temple 0.893 M 1.345 M (1.51$\uparrow$) 6929 7525 (1.09$\uparrow$) 13.03 15.26 (1.17$\uparrow$) 216 149 (1.45$\downarrow$)
TT-Train 1.113 M 1.332 M (1.20$\uparrow$) 6551 7128 (1.09$\uparrow$) 12.64 13.62 (1.08$\uparrow$) 193 163 (1.18$\downarrow$)
TT-Truck 2.533 M 3.630 M (1.43$\uparrow$) 8662 11014 (1.27$\uparrow$) 14.64 20.12 (1.37$\uparrow$) 156 88 (1.77$\downarrow$)
Table 4: Full unconstrained attack results ($\epsilon=\infty$) on all scenes from the NeRF-Synthetic, MIP-NeRF360 and Tanks-and-Temples datasets. All results are reported by averaging over three individual runs.
Unconstrained Poison-splat attack with $\epsilon=\infty$
Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes] Render speed [FPS]
Scene Setting clean poisoned clean poisoned clean poisoned clean poisoned
NS-chair 0.494 M 4.152 M (8.40$\uparrow$) 3752 51424 (13.71$\uparrow$) 9.45 40.71 (4.31$\uparrow$) 260 27 (9.63$\downarrow$)
NS-drums 0.385 M 4.109 M (10.67$\uparrow$) 3456 62328 (18.03$\uparrow$) 8.75 43.25 (4.94$\uparrow$) 347 24 (14.46$\downarrow$)
NS-ficus 0.265 M 3.718 M (14.03$\uparrow$) 3194 50450 (15.80$\uparrow$) 6.99 36.92 (5.28$\uparrow$) 346 31 (11.16$\downarrow$)
NS-hotdog 0.185 M 4.272 M (23.09$\uparrow$) 3336 47859 (14.35$\uparrow$) 8.57 38.85 (4.53$\uparrow$) 443 29 (15.28$\downarrow$)
NS-lego 0.341 M 4.159 M (12.20$\uparrow$) 3532 78852 (22.33$\uparrow$) 8.62 42.46 (4.93$\uparrow$) 349 25 (13.96$\downarrow$)
NS-materials 0.169 M 3.380 M (20.00$\uparrow$) 3172 43998 (13.87$\uparrow$) 7.33 32.40 (4.42$\uparrow$) 447 37 (12.08$\downarrow$)
NS-mic 0.205 M 3.940 M (19.22$\uparrow$) 3499 61835 (17.67$\uparrow$) 8.08 39.02 (4.83$\uparrow$) 300 29 (10.34$\downarrow$)
NS-ship 0.272 M 4.317 M (15.87$\uparrow$) 3692 80956 (21.93$\uparrow$) 8.87 44.11 (4.97$\uparrow$) 326 24 (13.58$\downarrow$)
MIP-bicycle 5.793 M 25.268 M (4.36$\uparrow$) 17748 63236 (3.56$\uparrow$) 33.42 81.48 (2.44$\uparrow$) 69 16 (4.31$\downarrow$)
MIP-bonsai 1.294 M 20.127 M (15.55$\uparrow$) 10216 54506 (5.34$\uparrow$) 21.68 75.18 (3.47$\uparrow$) 214 18 (11.89$\downarrow$)
MIP-counter 1.195 M 11.167 M (9.34$\uparrow$) 10750 80732 (7.51$\uparrow$) 26.46 62.04 (2.34$\uparrow$) 164 19 (8.63$\downarrow$)
MIP-flowers 3.508 M 18.075 M (5.15$\uparrow$) 12520 45515 (3.64$\uparrow$) 25.47 62.62 (2.46$\uparrow$) 124 24 (5.17$\downarrow$)
MIP-garden 5.684 M 21.527 M (3.79$\uparrow$) 17561 52140 (2.97$\uparrow$) 33.19 83.81 (2.53$\uparrow$) 81 17 (4.76$\downarrow$)
MIP-kitchen 1.799 M 12.830 M (7.13$\uparrow$) 12442 77141 (6.20$\uparrow$) 26.61 73.04 (2.74$\uparrow$) 133 16 (8.31$\downarrow$)
MIP-room 1.513 M 16.019 M (10.59$\uparrow$) 12316 57540 (4.67$\uparrow$) 25.36 76.25 (3.01$\uparrow$) 154 17 (9.06$\downarrow$)
MIP-stump 4.671 M 13.550 M (2.90$\uparrow$) 14135 36181 (2.56$\uparrow$) 27.06 51.51 (1.90$\uparrow$) 111 27 (4.11$\downarrow$)
MIP-treehill 3.398 M 13.634 M (4.01$\uparrow$) 11444 36299 (3.17$\uparrow$) 22.76 52.83 (2.32$\uparrow$) 123 24 (5.12$\downarrow$)
TT-Auditorium 0.699 M 4.153 M (5.94$\uparrow$) 5982 12666 (2.12$\uparrow$) 12.69 22.50 (1.77$\uparrow$) 266 76 (3.50$\downarrow$)
TT-Ballroom 3.103 M 7.534 M (2.43$\uparrow$) 10416 19832 (1.90$\uparrow$) 20.37 33.63 (1.65$\uparrow$) 96 46 (2.09$\downarrow$)
TT-Barn 1.005 M 5.456 M (5.43$\uparrow$) 7813 15682 (2.01$\uparrow$) 13.10 26.10 (1.99$\uparrow$) 229 61 (3.75$\downarrow$)
TT-Caterpillar 1.290 M 6.980 M (5.41$\uparrow$) 6624 18909 (2.85$\uparrow$) 11.96 29.51 (2.47$\uparrow$) 228 53 (4.30$\downarrow$)
TT-Church 2.351 M 7.564 M (3.22$\uparrow$) 10076 21296 (2.11$\uparrow$) 17.43 33.39 (1.92$\uparrow$) 132 43 (3.07$\downarrow$)
TT-Courthouse 0.604 M 3.388 M (5.61$\uparrow$) 11402 29856 (2.62$\uparrow$) 13.81 25.33 (1.83$\uparrow$) 284 54 (5.26$\downarrow$)
TT-Courtroom 2.890 M 13.196 M (4.57$\uparrow$) 9896 33871 (3.42$\uparrow$) 16.88 41.69 (2.47$\uparrow$) 139 41 (3.39$\downarrow$)
TT-Family 2.145 M 11.700 M (5.45$\uparrow$) 7261 27533 (3.79$\uparrow$) 13.58 38.79 (2.86$\uparrow$) 182 43 (4.23$\downarrow$)
TT-Francis 0.759 M 7.435 M (9.80$\uparrow$) 5208 19777 (3.80$\uparrow$) 10.19 31.01 (3.04$\uparrow$) 279 51 (5.47$\downarrow$)
TT-Horse 1.308 M 9.358 M (7.15$\uparrow$) 5125 22756 (4.44$\uparrow$) 12.05 34.90 (2.90$\uparrow$) 220 46 (4.78$\downarrow$)
TT-Ignatius 3.144 M 11.278 M (3.59$\uparrow$) 9943 28895 (2.91$\uparrow$) 16.32 39.61 (2.43$\uparrow$) 149 42 (3.55$\downarrow$)
TT-Lighthouse 0.888 M 5.081 M (5.72$\uparrow$) 7466 23842 (3.19$\uparrow$) 14.21 30.67 (2.16$\uparrow$) 213 47 (4.53$\downarrow$)
TT-M60 1.667 M 7.076 M (4.24$\uparrow$) 7948 21062 (2.65$\uparrow$) 14.08 31.82 (2.26$\uparrow$) 200 46 (4.35$\downarrow$)
TT-Meetingroom 1.267 M 7.066 M (5.58$\uparrow$) 6645 19182 (2.89$\uparrow$) 13.53 31.44 (2.32$\uparrow$) 197 50 (3.94$\downarrow$)
TT-Museum 4.439 M 16.501 M (3.72$\uparrow$) 12704 43317 (3.41$\uparrow$) 19.79 48.89 (2.47$\uparrow$) 121 36 (3.36$\downarrow$)
TT-Palace 0.711 M 2.764 M (3.89$\uparrow$) 8980 14065 (1.57$\uparrow$) 14.27 20.07 (1.41$\uparrow$) 185 84 (2.20$\downarrow$)
TT-Panther 1.812 M 8.112 M (4.48$\uparrow$) 8511 22638 (2.66$\uparrow$) 13.49 33.19 (2.46$\uparrow$) 199 49 (4.06$\downarrow$)
TT-Playground 2.309 M 10.306 M (4.46$\uparrow$) 8717 27304 (3.13$\uparrow$) 15.17 38.77 (2.56$\uparrow$) 151 39 (3.87$\downarrow$)
TT-Temple 0.893 M 5.231 M (5.86$\uparrow$) 6929 15238 (2.20$\uparrow$) 13.03 27.96 (2.15$\uparrow$) 216 52 (4.15$\downarrow$)
TT-Train 1.113 M 4.916 M (4.42$\uparrow$) 6551 14840 (2.27$\uparrow$) 12.64 25.34 (2.00$\uparrow$) 193 59 (3.27$\downarrow$)
TT-Truck 2.533 M 8.004 M (3.16$\uparrow$) 8662 21166 (2.44$\uparrow$) 14.64 33.26 (2.27$\uparrow$) 156 47 (3.32$\downarrow$)
C.2 Full black-box attack results
We report the full results of attacking a black-box victim (Scaffold-GS) (Lu et al., 2024) on NeRF-Synthetic and MIP-NeRF360 datasets in Table 5.
C.3 Ablation study on constraint strength and attack performance
Typically, attacker can choose to balance the attack performance and stealthiness by adjusting attack constraints $\epsilon$. To better understand the tradeoff between $\epsilon$ and attack performance, we poisoned each scene with three levels of $\epsilon$ setting, which are $8/255$, $16/255$ and $24/255$. The comparisons of computation cost under different constraint levels of attacks are provided in Table 6 (NeRF-Synthetic) and Table 7. All results have been reported via the average and standard deviation of three individual runs of victim training.
Table 5: Full black-box attack results on NeRF-Synthetic and MIP-NeRF360 datasets. The victim system, Scaffold-GS, utilizes distinctly different Gaussian Splatting feature representations compared to traditional Gaussian Splatting and remains unknown to the attacker. These results demonstrate the robust generalization ability of our attack against unknown black-box victim systems.
Constrained Black-box Poison-splat attack with $\epsilon=16/255$ against Scaffold-GS
Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes]
Scene Setting clean poisoned clean poisoned clean poisoned
NS-chair 1.000 M 1.615 M (1.61$\uparrow$) 3418 4003 (1.17$\uparrow$) 9.91 17.18 (1.73$\uparrow$)
NS-drums 1.000 M 1.771 M (1.77$\uparrow$) 4022 4382 (1.09$\uparrow$) 10.62 16.29 (1.53$\uparrow$)
NS-ficus 1.000 M 2.301 M (2.30$\uparrow$) 3416 4234 (1.24$\uparrow$) 9.64 18.54 (1.92$\uparrow$)
NS-hotdog 1.000 M 1.463 M (1.46$\uparrow$) 3539 4516 (1.28$\uparrow$) 10.74 18.62 (1.73$\uparrow$)
NS-lego 0.414 M 2.074 M (5.01$\uparrow$) 3003 4808 (1.60$\uparrow$) 9.77 17.91 (1.83$\uparrow$)
NS-materials 1.000 M 2.520 M (2.52$\uparrow$) 4110 5190 (1.26$\uparrow$) 10.87 17.75 (1.63$\uparrow$)
NS-mic 1.000 M 2.053 M (2.05$\uparrow$) 3682 5054 (1.37$\uparrow$) 10.36 18.53 (1.79$\uparrow$)
NS-ship 1.000 M 3.291 M (3.29$\uparrow$) 3492 5024 (1.44$\uparrow$) 11.68 21.92 (1.88$\uparrow$)
MIP-bicycle 8.883 M 16.947 M (1.91$\uparrow$) 12817 13650 (1.06$\uparrow$) 38.93 44.98 (1.16$\uparrow$)
MIP-bonsai 4.368 M 10.608 M (2.43$\uparrow$) 10080 12218 (1.21$\uparrow$) 35.33 37.71 (1.07$\uparrow$)
MIP-counter 2.910 M 6.387 M (2.19$\uparrow$) 12580 17702 (1.41$\uparrow$) 38.53 44.83 (1.16$\uparrow$)
MIP-flowers 7.104 M 11.614 M (1.63$\uparrow$) 8241 9766 (1.19$\uparrow$) 36.56 38.31 (1.05$\uparrow$)
MIP-garden 7.630 M 11.555 M (1.51$\uparrow$) 9194 10966 (1.19$\uparrow$) 38.57 42.79 (1.11$\uparrow$)
MIP-kitchen 3.390 M 6.698 M (1.98$\uparrow$) 14037 16895 (1.20$\uparrow$) 43.83 48.10 (1.10$\uparrow$)
MIP-room 2.846 M 10.527 M (3.70$\uparrow$) 13323 13981 (1.05$\uparrow$) 36.24 48.01 (1.32$\uparrow$)
MIP-stump 6.798 M 14.544 M (2.14$\uparrow$) 7322 9432 (1.29$\uparrow$) 33.53 36.32 (1.08$\uparrow$)
Unconstrained Black-box Poison-splat attack with $\epsilon=\infty$ against Scaffold-GS
Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes]
Scene Setting clean poisoned clean poisoned clean poisoned
NS-chair 1.000 M 3.082 M (3.08$\uparrow$) 3418 5197 (1.52$\uparrow$) 9.91 25.34 (2.56$\uparrow$)
NS-drums 1.000 M 5.026 M (5.03$\uparrow$) 4022 7364 (1.83$\uparrow$) 10.62 28.24 (2.66$\uparrow$)
NS-ficus 1.000 M 2.106 M (2.11$\uparrow$) 3416 4446 (1.30$\uparrow$) 9.64 21.65 (2.25$\uparrow$)
NS-hotdog 1.000 M 3.541 M (3.54$\uparrow$) 3539 5492 (1.55$\uparrow$) 10.74 24.93 (2.32$\uparrow$)
NS-lego 0.414 M 3.973 M (9.60$\uparrow$) 3003 6242 (2.08$\uparrow$) 9.77 26.11 (2.67$\uparrow$)
NS-materials 1.000 M 3.997 M (4.00$\uparrow$) 4110 5997 (1.46$\uparrow$) 10.87 25.48 (2.34$\uparrow$)
NS-mic 1.000 M 3.021 M (3.02$\uparrow$) 3682 5490 (1.49$\uparrow$) 10.36 23.71 (2.29$\uparrow$)
NS-ship 1.000 M 4.717 M (4.72$\uparrow$) 3492 6802 (1.95$\uparrow$) 11.68 28.22 (2.42$\uparrow$)
MIP-bicycle 8.883 M 33.284 M (3.75$\uparrow$) 12817 22042 (1.72$\uparrow$) 38.93 84.71 (2.18$\uparrow$)
MIP-bonsai 4.368 M 28.042 M (6.42$\uparrow$) 10080 22115 (2.19$\uparrow$) 35.33 78.36 (2.22$\uparrow$)
MIP-counter 2.910 M 12.928 M (4.44$\uparrow$) 12580 16168 (1.29$\uparrow$) 38.53 55.59 (1.44$\uparrow$)
MIP-flowers 7.104 M 27.610 M (3.89$\uparrow$) 8241 18352 (2.23$\uparrow$) 36.56 73.43 (2.01$\uparrow$)
MIP-garden 7.630 M 23.828 M (3.12$\uparrow$) 9194 20400 (2.22$\uparrow$) 38.57 85.45 (2.22$\uparrow$)
MIP-kitchen 3.390 M 14.404 M (4.25$\uparrow$) 14037 17838 (1.27$\uparrow$) 43.83 63.32 (1.44$\uparrow$)
MIP-room 2.846 M 21.060 M (7.40$\uparrow$) 13323 21672 (1.63$\uparrow$) 36.24 76.94 (2.12$\uparrow$)
MIP-stump 6.798 M 34.027 M (5.01$\uparrow$) 7322 20797 (2.84$\uparrow$) 33.53 79.64 (2.38$\uparrow$)
Table 6: Attack performance under different levels of constraints on NeRF Synthetic dataset. Standard deviation are reported over three individual runs.
NeRF-Synthetic Dataset
Scene Attack setting Number of Gaussians Peak GPU memory [MB] Training time [minutes]
chair clean 0.494 M $\pm$ 0.001 M 3752 MB $\pm$ 57 MB 9.45 min $\pm$ 0.55 min
$\epsilon=8/255$ 0.670 M $\pm$ 0.002 M 4520 MB $\pm$ 4 MB 11.24 min $\pm$ 0.09 min
$\epsilon=16/255$ 0.957 M $\pm$ 0.003 M 8003 MB $\pm$ 138 MB 13.41 min $\pm$ 0.44 min
$\epsilon=24/255$ 1.253 M $\pm$ 0.005 M 16394 MB $\pm$ 284 MB 15.89 min $\pm$ 0.68 min
unconstrained 4.152 M $\pm$ 0.006 M 51424 MB $\pm$ 98 MB 40.71 min $\pm$ 0.67 min
drums clean 0.385 M $\pm$ 0.001 M 3456 MB $\pm$ 47 MB 8.75 min $\pm$ 0.39 min
$\epsilon=8/255$ 0.469 M $\pm$ 0.000 M 4353 MB $\pm$ 53 MB 10.19 min $\pm$ 0.42 min
$\epsilon=16/255$ 0.695 M $\pm$ 0.002 M 7190 MB $\pm$ 261 MB 11.94 min $\pm$ 0.45 min
$\epsilon=24/255$ 0.966 M $\pm$ 0.007 M 14668 MB $\pm$ 451 MB 14.86 min $\pm$ 0.33 min
unconstrained 4.109 M $\pm$ 0.039 M 62328 MB $\pm$ 878 MB 43.25 min $\pm$ 0.75 min
ficus clean 0.265 M $\pm$ 0.000 M 3194 MB $\pm$ 42 MB 6.99 min $\pm$ 0.11 min
$\epsilon=8/255$ 0.246 M $\pm$ 0.001 M 3238 MB $\pm$ 1 MB 8.28 min $\pm$ 0.56 min
$\epsilon=16/255$ 0.273 M $\pm$ 0.001 M 3783 MB $\pm$ 51 MB 8.94 min $\pm$ 0.40 min
$\epsilon=24/255$ 0.450 M $\pm$ 0.001 M 8993 MB $\pm$ 112 MB 11.32 min $\pm$ 0.31 min
unconstrained 3.718 M $\pm$ 0.001 M 50450 MB $\pm$ 452 MB 36.92 min $\pm$ 0.86 min
hotdog clean 0.185 M $\pm$ 0.000 M 3336 MB $\pm$ 11 MB 8.57 min $\pm$ 0.30 min
$\epsilon=8/255$ 0.610 M $\pm$ 0.000 M 10362 MB $\pm$ 280 MB 12.96 min $\pm$ 0.23 min
$\epsilon=16/255$ 1.147 M $\pm$ 0.003 M 29747 MB $\pm$ 57 MB 17.43 min $\pm$ 0.61 min
$\epsilon=24/255$ 1.553 M $\pm$ 0.005 M 39087 MB $\pm$ 198 MB 21.26 min $\pm$ 0.99 min
unconstrained 4.272 M $\pm$ 0.003 M 47859 MB $\pm$ 688 MB 38.85 min $\pm$ 0.90 min
lego clean 0.341 M $\pm$ 0.001 M 3532 MB $\pm$ 91 MB 8.62 min $\pm$ 0.41 min
$\epsilon=8/255$ 0.469 M $\pm$ 0.002 M 4648 MB $\pm$ 120 MB 10.38 min $\pm$ 0.47 min
$\epsilon=16/255$ 0.805 M $\pm$ 0.001 M 9960 MB $\pm$ 132 MB 13.02 min $\pm$ 0.50 min
$\epsilon=24/255$ 1.116 M $\pm$ 0.011 M 18643 MB $\pm$ 504 MB 15.33 min $\pm$ 0.63 min
unconstrained 4.159 M $\pm$ 0.002 M 78852 MB $\pm$ 2392 MB 42.46 min $\pm$ 0.88 min
materials clean 0.169 M $\pm$ 0.001 M 3172 MB $\pm$ 32 MB 7.33 min $\pm$ 0.21 min
$\epsilon=8/255$ 0.229 M $\pm$ 0.000 M 3581 MB $\pm$ 52 MB 8.84 min $\pm$ 0.29 min
$\epsilon=16/255$ 0.410 M $\pm$ 0.001 M 4886 MB $\pm$ 66 MB 9.92 min $\pm$ 0.10 min
$\epsilon=24/255$ 0.589 M $\pm$ 0.000 M 7724 MB $\pm$ 106 MB 11.51 min $\pm$ 0.43 min
unconstrained 3.380 M $\pm$ 0.000 M 43998 MB $\pm$ 675 MB 32.40 min $\pm$ 1.02 min
mic clean 0.205 M $\pm$ 0.000 M 3499 MB $\pm$ 130 MB 8.08 min $\pm$ 0.20 min
$\epsilon=8/255$ 0.251 M $\pm$ 0.001 M 4038 MB $\pm$ 132 MB 9.29 min $\pm$ 0.03 min
$\epsilon=16/255$ 0.359 M $\pm$ 0.001 M 6437 MB $\pm$ 70 MB 10.96 min $\pm$ 0.10 min
$\epsilon=24/255$ 0.514 M $\pm$ 0.005 M 13014 MB $\pm$ 141 MB 12.89 min $\pm$ 0.52 min
unconstrained 3.940 M $\pm$ 0.004 M 61835 MB $\pm$ 586 MB 39.02 min $\pm$ 0.65 min
ship clean 0.272 M $\pm$ 0.001 M 3692 MB $\pm$ 52 MB 8.87 min $\pm$ 0.44 min
$\epsilon=8/255$ 0.516 M $\pm$ 0.003 M 5574 MB $\pm$ 55 MB 11.01 min $\pm$ 0.16 min
$\epsilon=16/255$ 1.071 M $\pm$ 0.003 M 16666 MB $\pm$ 548 MB 16.12 min $\pm$ 0.42 min
$\epsilon=24/255$ 1.365 M $\pm$ 0.002 M 29828 MB $\pm$ 192 MB 18.46 min $\pm$ 0.47 min
unconstrained 4.317 M $\pm$ 0.005 M 80956 MB $\pm$ 13 MB 44.11 min $\pm$ 0.94 min
Table 7: Attack performance under different levels of constraints on MIP-NeRF360 dataset. Standard deviation are reported over three individual runs.
MIP-NeRF360 Dataset
Scene Attack setting Number of Gaussians Peak GPU memory [MB] Training time [minutes]
bicycle clean 5.793 M $\pm$ 0.007 M 17748 MB $\pm$ 554 MB 33.42 min $\pm$ 0.22 min
$\epsilon=8/255$ 6.116 M $\pm$ 0.019 M 18608 MB $\pm$ 401 MB 34.73 min $\pm$ 0.67 min
$\epsilon=16/255$ 10.129 M $\pm$ 0.053 M 27074 MB $\pm$ 1180 MB 44.44 min $\pm$ 0.43 min
$\epsilon=24/255$ 13.265 M $\pm$ 0.080 M 34870 MB $\pm$ 1882 MB 51.37 min $\pm$ 0.26 min
unconstrained 25.268 M $\pm$ 0.627 M 63236 MB $\pm$ 537 MB 81.48 min $\pm$ 1.44 min
bonsai clean 1.294 M $\pm$ 0.002 M 10216 MB $\pm$ 21 MB 21.68 min $\pm$ 1.18 min
$\epsilon=8/255$ 1.778 M $\pm$ 0.011 M 11495 MB $\pm$ 264 MB 21.39 min $\pm$ 0.48 min
$\epsilon=16/255$ 6.150 M $\pm$ 0.063 M 20564 MB $\pm$ 58 MB 32.51 min $\pm$ 0.27 min
$\epsilon=24/255$ 9.321 M $\pm$ 0.058 M 27191 MB $\pm$ 194 MB 39.84 min $\pm$ 0.15 min
unconstrained 20.127 M $\pm$ 0.122 M 54506 MB $\pm$ 3437 MB 75.18 min $\pm$ 0.39 min
counter clean 1.195 M $\pm$ 0.005 M 10750 MB $\pm$ 104 MB 26.46 min $\pm$ 0.57 min
$\epsilon=8/255$ 1.739 M $\pm$ 0.010 M 15133 MB $\pm$ 408 MB 28.06 min $\pm$ 0.61 min
$\epsilon=16/255$ 4.628 M $\pm$ 0.014 M 29796 MB $\pm$ 187 MB 37.63 min $\pm$ 0.48 min
$\epsilon=24/255$ 6.649 M $\pm$ 0.002 M 47607 MB $\pm$ 796 MB 43.68 min $\pm$ 0.14 min
unconstrained 11.167 M $\pm$ 0.092 M 80732 MB $\pm$ 255 MB 62.04 min $\pm$ 0.13 min
flowers clean 3.508 M $\pm$ 0.008 M 12520 MB $\pm$ 263 MB 25.47 min $\pm$ 0.29 min
$\epsilon=8/255$ 3.457 M $\pm$ 0.026 M 12391 MB $\pm$ 262 MB 25.53 min $\pm$ 0.39 min
$\epsilon=16/255$ 5.177 M $\pm$ 0.008 M 16338 MB $\pm$ 566 MB 30.05 min $\pm$ 0.42 min
$\epsilon=24/255$ 7.280 M $\pm$ 0.129 M 20264 MB $\pm$ 44 MB 34.25 min $\pm$ 0.59 min
unconstrained 18.075 M $\pm$ 0.136 M 45515 MB $\pm$ 2810 MB 62.62 min $\pm$ 0.62 min
garden clean 5.684 M $\pm$ 0.014 M 17561 MB $\pm$ 194 MB 33.19 min $\pm$ 0.54 min
$\epsilon=8/255$ 5.122 M $\pm$ 0.006 M 16498 MB $\pm$ 108 MB 32.17 min $\pm$ 0.71 min
$\epsilon=16/255$ 7.697 M $\pm$ 0.016 M 23115 MB $\pm$ 505 MB 40.82 min $\pm$ 0.23 min
$\epsilon=24/255$ 10.356 M $\pm$ 0.010 M 28130 MB $\pm$ 174 MB 48.88 min $\pm$ 0.15 min
unconstrained 21.527 M $\pm$ 0.027 M 52140 MB $\pm$ 1208 MB 83.81 min $\pm$ 0.07 min
kitchen clean 1.799 M $\pm$ 0.002 M 12442 MB $\pm$ 252 MB 26.61 min $\pm$ 0.28 min
$\epsilon=8/255$ 2.675 M $\pm$ 0.001 M 13760 MB $\pm$ 117 MB 31.39 min $\pm$ 0.66 min
$\epsilon=16/255$ 6.346 M $\pm$ 0.046 M 20902 MB $\pm$ 353 MB 43.24 min $\pm$ 0.59 min
$\epsilon=24/255$ 9.445 M $\pm$ 0.040 M 27646 MB $\pm$ 614 MB 53.16 min $\pm$ 0.32 min
unconstrained 12.830 M $\pm$ 0.059 M 77141 MB $\pm$ 1444 MB 73.04 min $\pm$ 0.28 min
room clean 1.513 M $\pm$ 0.004 M 12316 MB $\pm$ 13 MB 25.36 min $\pm$ 0.09 min
$\epsilon=8/255$ 3.034 M $\pm$ 0.003 M 19542 MB $\pm$ 363 MB 32.99 min $\pm$ 0.34 min
$\epsilon=16/255$ 7.129 M $\pm$ 0.020 M 46238 MB $\pm$ 802 MB 47.98 min $\pm$ 0.72 min
$\epsilon=24/255$ 9.724 M $\pm$ 0.053 M 80142 MB $\pm$ 249 MB 56.87 min $\pm$ 0.51 min
unconstrained 16.019 M $\pm$ 0.025 M 57540 MB $\pm$ 3983 MB 76.25 min $\pm$ 0.14 min
stump clean 4.671 M $\pm$ 0.070 M 14135 MB $\pm$ 218 MB 27.06 min $\pm$ 0.31 min
$\epsilon=8/255$ 5.152 M $\pm$ 0.072 M 15342 MB $\pm$ 121 MB 28.21 min $\pm$ 0.31 min
$\epsilon=16/255$ 10.003 M $\pm$ 0.292 M 25714 MB $\pm$ 1806 MB 38.48 min $\pm$ 0.59 min
$\epsilon=24/255$ 14.292 M $\pm$ 0.050 M 34246 MB $\pm$ 1108 MB 46.89 min $\pm$ 0.49 min
unconstrained 13.550 M $\pm$ 1.257 M 36181 MB $\pm$ 3769 MB 51.51 min $\pm$ 1.90 min
Appendix D Notations
We provide Table 8 to list all the notations appeared in the paper.
Table 8: Notation List
Variable Description
$\mathcal{G}$ 3D Gaussians
$\|\mathcal{G}\|$ The number of 3D Gaussians
$G_{i}$ The $i$-th Gaussian in the scene
$\mathcal{D}$ Clean image datasets
$\mathcal{D}_{p}$ Poisoned image datasets
$V_{k}$ The $k$-th camera view’s clean image
$P_{k}$ The $k$-th camera view’s camera pose
$\tilde{V_{k}}$ The $k$-th camera view’s poisoned image
$\tilde{P_{k}}$ The $k$-th camera view’s poisoned pose
${V^{\prime}_{k}}$ The $k$-th camera view’s rendered image
$N$ The total number of views
$\mu\in\mathbb{R}^{3}$ Center 3D coordinates (positions) of a Gaussian
$s\in\mathbb{R}^{3}$ 3D scales of a Gaussian
$\alpha\in\mathbb{R}$ Opacity of a Gaussian
$q\in\mathbb{R}^{4}$ Rotation quaternion of a Gaussian
$c\in\mathbb{R}^{3\times(d+1)^{2}}$ Color of a Gaussian represented as spherical harmonics, where $d$ is the degree
$\tau_{g}$ Gradient threshold
$\tau_{s}$ Scale threshold
$\tau_{\alpha}$ Opacity threshold
$\mathcal{R}$ Renderer function
$\epsilon$ Perturbation range
$\eta$ Perturbation step size
$T$ The iteration number of inner optimization
$\tilde{T}$ The iteration number of outer optimization
$\mathcal{P}_{\epsilon}(\cdot)$ Projection into $\epsilon$-ball
$\mathcal{L}$ Reconstruction loss
$\mathcal{L}_{1}$ $\mathcal{L}_{1}$ loss
$\mathcal{L}_{\text{D-SSIM}}$ Structural similarity index measure (SSIM) loss
$\mathcal{C}(\mathcal{G}^{*})$ Training costs of 3D Gaussians
$\mathcal{S}_{\text{TV}}(V_{k})$ Total variation score
$\lambda$ Loss trade-off parameters
Appendix E Visualizations
In this section, we present both the attacker-crafted poisoned images and the resultant reconstructions by the victim model. The visualizations aim to provide readers with more straightforward understanding of our attack. The constrained attack visualizations are provided in Table 9, and the unconstrained attack visualizations are provided in Table 10.
We have following interesting observations:
  • Unconstrained Poisoned Attacks: On both NeRF-Synthetic and MIP-NeRF360, the reconstruction results consistently show a low PSNR (around 20 dB). While the poisoned images and victim reconstructions may both resemble chaotic arrangements of lines with needle-like Gaussians, subtle differences in detail exist. These differences, though difficult for the human eye to distinguish, are reflected in the PSNR.
  • Constrained Attacks on NeRF-Synthetic: the reconstruction results are significantly better than unconstrained attacks, with an average PSNR of around 30 dB.
  • Constrained Attacks on MIP-NeRF360: There are substantial appearance differences between the victim reconstructions and the poisoned input images, which are easily noticeable through human observation. Specifically, in the bicycle scene, the reconstructed trees appear to have denser leaves. In the bonsai and counter scenes, we observe notable changes in reflective areas, with several reflection highlights disappearing in the victim’s reconstructed images. Additionally, in the counter scene, there is evidence of underfitting, where the complex textures of the poisoned data degrade into large, blurry floaters in the victim’s reconstruction.
Table 9: Attacker poisoned and victim reconstructed images with PSNR values under constrained attack ($\epsilon=16/255$) settings.
Dataset Setting Attacker Poisoned Image Victim Reconstructed Image PSNR
NS-Chair [Uncaptioned image] [Uncaptioned image] 37.07 dB
NS-Drums [Uncaptioned image] [Uncaptioned image] 30.32 dB
NS-Ficus [Uncaptioned image] [Uncaptioned image] 35.77 dB
MIP-bicycle [Uncaptioned image] [Uncaptioned image] 18.20 dB
MIP-bonsai [Uncaptioned image] [Uncaptioned image] 22.67 dB
MIP-counter [Uncaptioned image] [Uncaptioned image] 24.45 dB
Table 10: Attacker poisoned and victim reconstructed images with PSNR values under unconstrained attack settings.
Dataset Scene Attacker Poisoned Image Victim Reconstructed Image PSNR
NS-Chair [Uncaptioned image] [Uncaptioned image] 19.54 dB
NS-Drums [Uncaptioned image] [Uncaptioned image] 18.65 dB
NS-Ficus [Uncaptioned image] [Uncaptioned image] 21.00 dB
MIP-bicycle [Uncaptioned image] [Uncaptioned image] 20.57 dB
MIP-bonsai [Uncaptioned image] [Uncaptioned image] 23.66 dB
MIP-counter [Uncaptioned image] [Uncaptioned image] 21.33 dB

### v2

@@
 These strategies not only ensure the effectiveness of our attack but also make it difficult to defend with simple defensive measures.
-We hope the revelation of this novel attack surface can spark attention to this crucial yet overlooked vulnerability of 3DGS systems [^1] Our code is available at https://github.com/jiahaolu97/poison-splat..
+We hope the revelation of this novel attack surface can spark attention to this crucial yet overlooked vulnerability of 3DGS systems.
+[^1] Our code is available at https://github.com/jiahaolu97/poison-splat.
 1 Introduction
@@
 In this paper, we reveal a severe yet unnoticed attack vector: the flexibility of 3DGS complexity can be abused to over-consume computation resources such as GPU memory and significantly drag down training speed of Gaussian Splatting system, pushing the training process to its worst computational complexity.
-We develop a computation cost attack method named Poison-splat as a proof-of-concept of this novel attack vector.
+We develop a computation cost attack method Poison-splat as a proof-of-concept of this novel attack vector.
 Poison-splat takes the form of training data poisoning (Tian et al., 2022), where the attacker manipulates the input data fed into the victim 3DGS system.
@@
 An attacker can hide themselves among normal users to submit poisoned data and launch the attack stealthily, or even secretly manipulate data uploaded by other users.
-During peak usage periods, this attack competes for computing resources with legitimate users, slowing down service response time and potentially causing severe consequences like service crashes, leading to financial losses for service providers.
+During peak usage periods, this attack competes for computing resources with legitimate users, slowing down service response times and potentially causing severe consequences like service crashes, leading to significant financial losses for service providers.
 We formulate Poison-splat attack as a max-min problem.
@@
 Following their original implementation, we use the recommended default hyper-parameters, which were proved effective across a broad spectrum of scenes from various datasets.
-For testing black-box attack effectiveness, we use the official implementation of Scaffold-GS [^6] https://github.com/city-super/Scaffold-GS (Lu et al., 2024) as a black box victim.
+For testing black-box attack effectiveness, we use the official implementation of Scaffold-GS [^6] https://github.com/city-super/Scaffold-GS (Lu et al., 2024b) as a black box victim.
 All experiments reported in the paper were carried out on a single NVIDIA A800-SXM4-80G GPU.
@@
 The full results containing all scenes are listed in the Appendix Table 3 and Table 4.
-We provide visualizations of example poisoned datasets and corresponding victim reconstructions in Appendix Section E.
+We provide visualizations of example poisoned datasets and corresponding victim reconstructions in Appendix Section F.
 We want to highlight some observations from the results.
@@
 The answer is affirmative.
-To test the effectiveness of our attack on a black-box victim, we chose Scaffold-GS (Lu et al., 2024), a 3DGS variant algorithm designed with a focus on efficiency.
+To test the effectiveness of our attack on a black-box victim, we chose Scaffold-GS (Lu et al., 2024b), a 3DGS variant algorithm designed with a focus on efficiency.
 Scaffold-GS differs significantly from traditional Gaussian Splatting as it utilizes anchor points to distribute local 3D Gaussians, which greatly alters the underlying feature representation.
@@
 This focus could significantly enhance the security and reliability of 3DGS systems in practical applications.
-Societal Impacts.
+Ethics Statement.
 While our methods could potentially be misused by malicious actors to disrupt 3DGS service providers and cause financial harm, our aim is not to facilitate such actions.
@@
 6 Acknowledgement
+This project is supported in part by NUS Start-up Grant A-0010106-00-00, and by the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity R&D Programme and CyberSG R&D Cyber Research Programme Office (Award: CRPO-GC1-NTU-002).
 We sincerely thank Xuanyu Yi (NTU), Xinjie Zhang (HKUST), Jiaqi Ma (WHU) and Shizun Wang (NUS) for their helpful discussions.
@@
 The Appendix is organized as follows: Section A discusses related works; Section B provides a formal and comprehensive description of 3D Gaussian Splatting training.
-Section C offers the full results of experiments mentioned in the main paper, presented here due to page constraints.
-Section D lists all the notations used throughout the paper.
-Section E presents examples of attacker-crafted poisoned images and the corresponding reconstructions from victim models trained with those images.
+Section C explicitly describes the threat model in this paper, aiming to provide readers with a clearer understanding.
+Section D offers the full results of experiments mentioned in the main paper, presented here due to page constraints.
+Section E lists all the notations used throughout the paper.
+Section F presents examples of attacker-crafted poisoned images and the corresponding reconstructions from victim models trained with those images.
+Section G reports the attack performance under various levels of perturbation intensity, ranging from $\epsilon=8/255$ to $\epsilon=24/255$.
+Section H studies how effective the Poison-Splat attack will be if the poison ratio is not 100%.
+Section I shows the defensive threshold of limiting Gaussian numbers and discuss the difficulty of setting a uniform defense threshold.
+Section J discusses another potential defense: image smoothing as a pre-processing, and shows that it is not an ideal defense in terms of keeping reconstruction quality.
+Section K reports and analyzes the time complexity of Poison-Splat attack.
 Appendix A Related Works
@@
 The explicit nature of 3D Gaussian Splatting enables real-time rendering capabilities and unparalleled levels of control and editability, demonstrating its effectiveness in 3D generation.
-Since its introduction in (Kerbl et al., 2023), 3D Gaussian Splatting has facilitated a diverse array of downstream applications, including text-to-3D generation (Chen et al., 2024; Tang et al., 2024; Yi et al., 2024), Simultaneous Localization and Mapping (Keetha et al., 2024; Matsuki et al., 2024; Yan et al., 2024), human avatars (Kocabas et al., 2024; Lei et al., 2024; Zielonka et al., 2023), 4D-dynamic scene modeling (Luiten et al., 2024; Wu et al., 2024a; Yang et al., 2024), and other applications (Xie et al., 2024b; Shen et al., 2024; Ye et al., 2023; Wang et al., 2024b; a).
+Since its introduction in (Kerbl et al., 2023), 3D Gaussian Splatting has facilitated a diverse array of downstream applications, including text-to-3D generation (Chen et al., 2024; Tang et al., 2024; Yi et al., 2024), Simultaneous Localization and Mapping (Keetha et al., 2024; Matsuki et al., 2024; Yan et al., 2024a), human avatars (Kocabas et al., 2024; Lei et al., 2024; Zielonka et al., 2023), 4D-dynamic scene modeling (Luiten et al., 2024; Wu et al., 2024a; Yang et al., 2024), and other applications (Xie et al., 2024b; Shen et al., 2024; Ye et al., 2023; Wang et al., 2024b; a).
 While the majority of the audience in the 3D vision community places their focus on utility, we found only a handful of research works put efforts into investigating the security issues of 3D algorithms.
@@
 A.2 Attacks on Machine Learning Systems
-Machine learning has been shown to be vulnerable to various types of attacks (Pitropakis et al., 2019; Rigaki & Garcia, 2023; Chakraborty et al., 2018; Oliynyk et al., 2023; Tian et al., 2022).
-The majority of attacks target the either confidentiality (including membership inference (Shokri et al., 2017; Salem et al., 2019; Choquette-Choo et al., 2021), data reconstruction attacks (Fredrikson et al., 2015; Carlini et al., 2021; 2023; Yu et al., 2023; Geiping et al., 2020; Lu et al., 2022) and model stealing attacks (Tramèr et al., 2016; Chandrasekaran et al., 2020)) or integrity (like adversarial attacks (Biggio et al., 2013; Szegedy et al., 2013; Dong et al., 2018; 2019) and data poisoning attacks (Barreno et al., 2010; Jagielski et al., 2018; Biggio et al., 2012; Mei & Zhu, 2015b)).
+Machine learning has been shown to be vulnerable to various types of attacks (Pitropakis et al., 2019; Rigaki & Garcia, 2023; Chakraborty et al., 2018; Oliynyk et al., 2023; Tian et al., 2022; Zhang et al., 2021; 2023).
+The majority of attacks target the either confidentiality (including membership inference (Shokri et al., 2017; Salem et al., 2019; Choquette-Choo et al., 2021), data reconstruction attacks (Fredrikson et al., 2015; Carlini et al., 2021; 2023; Yu et al., 2023; Geiping et al., 2020; Lu et al., 2022) and model stealing attacks (Tramèr et al., 2016; Chandrasekaran et al., 2020)) or integrity (like adversarial attacks (Biggio et al., 2013; Szegedy et al., 2013; Dong et al., 2018; 2019; Lu et al., 2024a; Zhou et al., ) and data poisoning attacks (Barreno et al., 2010; Jagielski et al., 2018; Biggio et al., 2012; Mei & Zhu, 2015b)).
 Different from the above mentioned attacks, we aim to illustrate that by carefully crafting input images, the attacker can indeed attack availability, i.e., timely and cost-affordable access to machine learning service.
@@
 Different from the work (Shumailov et al., 2021), our attack targets 3D reconstruction algorithms and is not driven by neural networks.
-A.2.1 Data Poisoning Attack
+A.2.1 Data Poisoning Attacks
 We consider our proposed attacks to fall into the category of data poisoning attack, which manipulates training data to launch the attack at the training stage.
@@
 For many machine learning systems, given the inherent complexity and their reliance on large datasets, it is intractable to exactly do the bi-level optimization.
-Muñoz-González et al.
-(2017) exploited back-gradient optimization technique to approximate real gradients through reverse-mode differentiation.
-Huang et al.
-(2020) formulated as a meta-learning problem, and approximated the inner optimal solution by training only $K$ SGD steps for each outer objective evaluation.
-Koh et al.
-proposed three methods to approximate the expensive bi-level optimization: influence functions (Koh & Liang, 2017), minimax duality and Karush-Kuhn-Tucker (KKT) conditions.
-Fowl et al.
-(2021a) used gradient matching which assumes inner optimality has been achieved by a fixed model.
-Fowl et al.
-(2021b) avoided the bi-level optimization with a well-trained proxy model from clean data.
+The work (Muñoz-González et al., 2017) exploited back-gradient optimization technique to approximate real gradients through reverse-mode differentiation.
+The work (Huang et al., 2020) formulated as a meta-learning problem, and approximated the inner optimal solution by training only $K$ SGD steps for each outer objective evaluation.
+The work (Koh et al., 2022) proposed three methods to approximate the expensive bi-level optimization: influence functions (Koh & Liang, 2017), minimax duality and Karush-Kuhn-Tucker (KKT) conditions.
+The work (Fowl et al., 2021a) used gradient matching which assumes inner optimality has been achieved by a fixed model.
+The work (Fowl et al., 2021b) avoided the bi-level optimization with a well-trained proxy model from clean data.
 Different from these previous studies, we propose a gradient-free optimization process to solve the outer optimization pipeline, and show a good optimization result on the attacker’s goal.
@@
 The adaptive density control is performed during regular optimization intervals and follows specific, predefined rules.
-In the initial design (Kerbl et al., 2023), for a Gaussian $G_{i}$ with view-space projection point $\mu^{k}_{i}=(\mu^{k}_{i,x},\mu^{k}_{i,y})$ under viewpoint $V_{k}$, the average view-space positional gradient $\nabla_{\mu_{i}}\mathcal{L}$ is calculated every 100 training iterations:
+In the initial design (Kerbl et al., 2023), for a Gaussian $G_{i}$ with view-space projection point $\mu^{k}_{i}=(\mu^{k}_{i,x},\mu^{k}_{i,y})$ under viewpoint $V_{k}$, the average view-space positional gradient $\nabla_{\mu_{i}}\mathcal{L}$ is calculated every 100 training iterations (Ye et al., 2024):
 $\nabla_{\mu_{i}}\mathcal{L}=\frac{\sum_{k=1}^{M}\|\frac{\partial\mathcal{L}(V^{\prime}_{k},{V}_{k})}{\partial\mu_{i}^{k}}\|}{M},$ (7)
@@
 We also notice some recent 3DGS studies (Cheng et al., 2024; Fang & Wang, 2024; Li et al., 2024; Bulò et al., 2024) that adopt different yet still pre-defined densification rules.
-Appendix C Extended Experiments
-C.1 Full attack results
+Appendix C Threat Model
+We provide a structured description of the attack’s goal, input, output, information and constraints in this section, to provide more clarity of the threat model:
+Attacker Goal: To increase the computation cost and over-consume computation resource of 3DGS service provider.
+Attacker Input: Clean data.
+Attacker Output: Poisoned data by running attack on the clean data.
+Attacker Information: Attacker doesn’t need to know the underlying hardware device of the victim.
+  • White-box: Attacker knows the specific 3DGS configuration of the victim.
+  • Black-box: Attacker only knows victim is using 3DGS, but does not know which variant of 3DGS representation and what configuration the victim is using.
+Attacker Constraint: Attacker can optionally choose to constrain the perturbation range $\epsilon$ of his additive perturbation to add on clean data.
+Attacker Algorithm: Algorithm 1.
+Appendix D Extended Experiments
+D.1 Full attack results
 We report the constrained attack results with $\epsilon=16/255$ on all scenes of NeRF-Synthetic, Mip-NeRF360 and Tanks-and-Temples dataset in Table 3, and the unconstrained attack results in Table 4.
@@
 TT-Museum 4.439 M 7.130 M (1.61$\uparrow$) 12704 18790 (1.48$\uparrow$) 19.79 26.38 (1.33$\uparrow$) 121 78 (1.55$\downarrow$)
+TT-Palace 0.711 M 0.696 M (0.98$\uparrow$) 8980 8372 (0.93$\uparrow$) 14.27 14.54 (1.02$\uparrow$) 185 186 (0.99$\downarrow$)
 TT-Panther 1.812 M 3.457 M (1.91$\uparrow$) 8511 12048 (1.42$\uparrow$) 13.49 19.46 (1.44$\uparrow$) 199 103 (1.93$\downarrow$)
@@
 TT-Truck 2.533 M 8.004 M (3.16$\uparrow$) 8662 21166 (2.44$\uparrow$) 14.64 33.26 (2.27$\uparrow$) 156 47 (3.32$\downarrow$)
-C.2 Full black-box attack results
-We report the full results of attacking a black-box victim (Scaffold-GS) (Lu et al., 2024) on NeRF-Synthetic and MIP-NeRF360 datasets in Table 5.
-C.3 Ablation study on constraint strength and attack performance
-Typically, attacker can choose to balance the attack performance and stealthiness by adjusting attack constraints $\epsilon$.
-To better understand the tradeoff between $\epsilon$ and attack performance, we poisoned each scene with three levels of $\epsilon$ setting, which are $8/255$, $16/255$ and $24/255$.
-The comparisons of computation cost under different constraint levels of attacks are provided in Table 6 (NeRF-Synthetic) and Table 7.
-All results have been reported via the average and standard deviation of three individual runs of victim training.
+D.2 Full black-box attack results
+We report the full results of attacking a black-box victim (Scaffold-GS) (Lu et al., 2024b) on NeRF-Synthetic and MIP-NeRF360 datasets in Table 5.
+To further test the effectiveness of Poison-Splat, we conduct black-box attack experiments on Mip-Splatting (Yu et al., 2024), which received the best student paper award at CVPR 2024 and has been highly popular (with 178 citations as of November 20th, 2024).
+Mip-Splatting is an advanced variant of the original 3D Gaussian Splatting, incorporating a 3D smoothing filter and a 2D Mip filter.
+These enhancements help eliminate various artifacts and achieve alias-free renderings.
+Results are shown in Table 6.
+We found that Mip-Splatting (Yu et al., 2024) consumes more GPU memory compared to the original 3D Gaussian Splatting, making it more prone to the worst attack consequence of running Out-of-Memory.
+As illustrated in Table 6, even when the attack perturbation is constrained to $\epsilon=16/255$, the GPU memory consumption nearly reached the 80 GB capacity of Nvidia A800.
+When we apply an unconstrained attack, all scenes in the MIP-NeRF360 dataset will result in denial-of-service.
+We examined the code implementation of Mip-Splatting [^7] https://github.com/autonomousvision/mip-splatting, and identified that it uses quantile computation in its Gaussian model densification function  [^8] In Line 524 of mip-splatting/scene/gaussian_model.py which requires massive GPU memory and easily triggers out-of-memory.
+Our attack highlights the vulnerability in various 3DGS algorithm implementations.
 Table 5: Full black-box attack results on NeRF-Synthetic and MIP-NeRF360 datasets.
@@
 MIP-stump 6.798 M 34.027 M (5.01$\uparrow$) 7322 20797 (2.84$\uparrow$) 33.53 79.64 (2.38$\uparrow$)
-Table 6: Attack performance under different levels of constraints on NeRF Synthetic dataset. Standard deviation are reported over three individual runs.
-NeRF-Synthetic Dataset
-Scene Attack setting Number of Gaussians Peak GPU memory [MB] Training time [minutes]
-chair clean 0.494 M $\pm$ 0.001 M 3752 MB $\pm$ 57 MB 9.45 min $\pm$ 0.55 min
-$\epsilon=8/255$ 0.670 M $\pm$ 0.002 M 4520 MB $\pm$ 4 MB 11.24 min $\pm$ 0.09 min
-$\epsilon=16/255$ 0.957 M $\pm$ 0.003 M 8003 MB $\pm$ 138 MB 13.41 min $\pm$ 0.44 min
-$\epsilon=24/255$ 1.253 M $\pm$ 0.005 M 16394 MB $\pm$ 284 MB 15.89 min $\pm$ 0.68 min
-unconstrained 4.152 M $\pm$ 0.006 M 51424 MB $\pm$ 98 MB 40.71 min $\pm$ 0.67 min
-drums clean 0.385 M $\pm$ 0.001 M 3456 MB $\pm$ 47 MB 8.75 min $\pm$ 0.39 min
-$\epsilon=8/255$ 0.469 M $\pm$ 0.000 M 4353 MB $\pm$ 53 MB 10.19 min $\pm$ 0.42 min
-$\epsilon=16/255$ 0.695 M $\pm$ 0.002 M 7190 MB $\pm$ 261 MB 11.94 min $\pm$ 0.45 min
-$\epsilon=24/255$ 0.966 M $\pm$ 0.007 M 14668 MB $\pm$ 451 MB 14.86 min $\pm$ 0.33 min
-unconstrained 4.109 M $\pm$ 0.039 M 62328 MB $\pm$ 878 MB 43.25 min $\pm$ 0.75 min
-ficus clean 0.265 M $\pm$ 0.000 M 3194 MB $\pm$ 42 MB 6.99 min $\pm$ 0.11 min
-$\epsilon=8/255$ 0.246 M $\pm$ 0.001 M 3238 MB $\pm$ 1 MB 8.28 min $\pm$ 0.56 min
-$\epsilon=16/255$ 0.273 M $\pm$ 0.001 M 3783 MB $\pm$ 51 MB 8.94 min $\pm$ 0.40 min
-$\epsilon=24/255$ 0.450 M $\pm$ 0.001 M 8993 MB $\pm$ 112 MB 11.32 min $\pm$ 0.31 min
-unconstrained 3.718 M $\pm$ 0.001 M 50450 MB $\pm$ 452 MB 36.92 min $\pm$ 0.86 min
-hotdog clean 0.185 M $\pm$ 0.000 M 3336 MB $\pm$ 11 MB 8.57 min $\pm$ 0.30 min
-$\epsilon=8/255$ 0.610 M $\pm$ 0.000 M 10362 MB $\pm$ 280 MB 12.96 min $\pm$ 0.23 min
-$\epsilon=16/255$ 1.147 M $\pm$ 0.003 M 29747 MB $\pm$ 57 MB 17.43 min $\pm$ 0.61 min
-$\epsilon=24/255$ 1.553 M $\pm$ 0.005 M 39087 MB $\pm$ 198 MB 21.26 min $\pm$ 0.99 min
-unconstrained 4.272 M $\pm$ 0.003 M 47859 MB $\pm$ 688 MB 38.85 min $\pm$ 0.90 min
-lego clean 0.341 M $\pm$ 0.001 M 3532 MB $\pm$ 91 MB 8.62 min $\pm$ 0.41 min
-$\epsilon=8/255$ 0.469 M $\pm$ 0.002 M 4648 MB $\pm$ 120 MB 10.38 min $\pm$ 0.47 min
-$\epsilon=16/255$ 0.805 M $\pm$ 0.001 M 9960 MB $\pm$ 132 MB 13.02 min $\pm$ 0.50 min
-$\epsilon=24/255$ 1.116 M $\pm$ 0.011 M 18643 MB $\pm$ 504 MB 15.33 min $\pm$ 0.63 min
-unconstrained 4.159 M $\pm$ 0.002 M 78852 MB $\pm$ 2392 MB 42.46 min $\pm$ 0.88 min
-materials clean 0.169 M $\pm$ 0.001 M 3172 MB $\pm$ 32 MB 7.33 min $\pm$ 0.21 min
-$\epsilon=8/255$ 0.229 M $\pm$ 0.000 M 3581 MB $\pm$ 52 MB 8.84 min $\pm$ 0.29 min
-$\epsilon=16/255$ 0.410 M $\pm$ 0.001 M 4886 MB $\pm$ 66 MB 9.92 min $\pm$ 0.10 min
-$\epsilon=24/255$ 0.589 M $\pm$ 0.000 M 7724 MB $\pm$ 106 MB 11.51 min $\pm$ 0.43 min
-unconstrained 3.380 M $\pm$ 0.000 M 43998 MB $\pm$ 675 MB 32.40 min $\pm$ 1.02 min
-mic clean 0.205 M $\pm$ 0.000 M 3499 MB $\pm$ 130 MB 8.08 min $\pm$ 0.20 min
-$\epsilon=8/255$ 0.251 M $\pm$ 0.001 M 4038 MB $\pm$ 132 MB 9.29 min $\pm$ 0.03 min
-$\epsilon=16/255$ 0.359 M $\pm$ 0.001 M 6437 MB $\pm$ 70 MB 10.96 min $\pm$ 0.10 min
-$\epsilon=24/255$ 0.514 M $\pm$ 0.005 M 13014 MB $\pm$ 141 MB 12.89 min $\pm$ 0.52 min
-unconstrained 3.940 M $\pm$ 0.004 M 61835 MB $\pm$ 586 MB 39.02 min $\pm$ 0.65 min
-ship clean 0.272 M $\pm$ 0.001 M 3692 MB $\pm$ 52 MB 8.87 min $\pm$ 0.44 min
-$\epsilon=8/255$ 0.516 M $\pm$ 0.003 M 5574 MB $\pm$ 55 MB 11.01 min $\pm$ 0.16 min
-$\epsilon=16/255$ 1.071 M $\pm$ 0.003 M 16666 MB $\pm$ 548 MB 16.12 min $\pm$ 0.42 min
-$\epsilon=24/255$ 1.365 M $\pm$ 0.002 M 29828 MB $\pm$ 192 MB 18.46 min $\pm$ 0.47 min
-unconstrained 4.317 M $\pm$ 0.005 M 80956 MB $\pm$ 13 MB 44.11 min $\pm$ 0.94 min
-Table 7: Attack performance under different levels of constraints on MIP-NeRF360 dataset. Standard deviation are reported over three individual runs.
-MIP-NeRF360 Dataset
-Scene Attack setting Number of Gaussians Peak GPU memory [MB] Training time [minutes]
-bicycle clean 5.793 M $\pm$ 0.007 M 17748 MB $\pm$ 554 MB 33.42 min $\pm$ 0.22 min
-$\epsilon=8/255$ 6.116 M $\pm$ 0.019 M 18608 MB $\pm$ 401 MB 34.73 min $\pm$ 0.67 min
-$\epsilon=16/255$ 10.129 M $\pm$ 0.053 M 27074 MB $\pm$ 1180 MB 44.44 min $\pm$ 0.43 min
-$\epsilon=24/255$ 13.265 M $\pm$ 0.080 M 34870 MB $\pm$ 1882 MB 51.37 min $\pm$ 0.26 min
-unconstrained 25.268 M $\pm$ 0.627 M 63236 MB $\pm$ 537 MB 81.48 min $\pm$ 1.44 min
-bonsai clean 1.294 M $\pm$ 0.002 M 10216 MB $\pm$ 21 MB 21.68 min $\pm$ 1.18 min
-$\epsilon=8/255$ 1.778 M $\pm$ 0.011 M 11495 MB $\pm$ 264 MB 21.39 min $\pm$ 0.48 min
-$\epsilon=16/255$ 6.150 M $\pm$ 0.063 M 20564 MB $\pm$ 58 MB 32.51 min $\pm$ 0.27 min
-$\epsilon=24/255$ 9.321 M $\pm$ 0.058 M 27191 MB $\pm$ 194 MB 39.84 min $\pm$ 0.15 min
-unconstrained 20.127 M $\pm$ 0.122 M 54506 MB $\pm$ 3437 MB 75.18 min $\pm$ 0.39 min
-counter clean 1.195 M $\pm$ 0.005 M 10750 MB $\pm$ 104 MB 26.46 min $\pm$ 0.57 min
-$\epsilon=8/255$ 1.739 M $\pm$ 0.010 M 15133 MB $\pm$ 408 MB 28.06 min $\pm$ 0.61 min
-$\epsilon=16/255$ 4.628 M $\pm$ 0.014 M 29796 MB $\pm$ 187 MB 37.63 min $\pm$ 0.48 min
-$\epsilon=24/255$ 6.649 M $\pm$ 0.002 M 47607 MB $\pm$ 796 MB 43.68 min $\pm$ 0.14 min
-unconstrained 11.167 M $\pm$ 0.092 M 80732 MB $\pm$ 255 MB 62.04 min $\pm$ 0.13 min
-flowers clean 3.508 M $\pm$ 0.008 M 12520 MB $\pm$ 263 MB 25.47 min $\pm$ 0.29 min
-$\epsilon=8/255$ 3.457 M $\pm$ 0.026 M 12391 MB $\pm$ 262 MB 25.53 min $\pm$ 0.39 min
-$\epsilon=16/255$ 5.177 M $\pm$ 0.008 M 16338 MB $\pm$ 566 MB 30.05 min $\pm$ 0.42 min
-$\epsilon=24/255$ 7.280 M $\pm$ 0.129 M 20264 MB $\pm$ 44 MB 34.25 min $\pm$ 0.59 min
-unconstrained 18.075 M $\pm$ 0.136 M 45515 MB $\pm$ 2810 MB 62.62 min $\pm$ 0.62 min
-garden clean 5.684 M $\pm$ 0.014 M 17561 MB $\pm$ 194 MB 33.19 min $\pm$ 0.54 min
-$\epsilon=8/255$ 5.122 M $\pm$ 0.006 M 16498 MB $\pm$ 108 MB 32.17 min $\pm$ 0.71 min
-$\epsilon=16/255$ 7.697 M $\pm$ 0.016 M 23115 MB $\pm$ 505 MB 40.82 min $\pm$ 0.23 min
-$\epsilon=24/255$ 10.356 M $\pm$ 0.010 M 28130 MB $\pm$ 174 MB 48.88 min $\pm$ 0.15 min
-unconstrained 21.527 M $\pm$ 0.027 M 52140 MB $\pm$ 1208 MB 83.81 min $\pm$ 0.07 min
-kitchen clean 1.799 M $\pm$ 0.002 M 12442 MB $\pm$ 252 MB 26.61 min $\pm$ 0.28 min
-$\epsilon=8/255$ 2.675 M $\pm$ 0.001 M 13760 MB $\pm$ 117 MB 31.39 min $\pm$ 0.66 min
-$\epsilon=16/255$ 6.346 M $\pm$ 0.046 M 20902 MB $\pm$ 353 MB 43.24 min $\pm$ 0.59 min
-$\epsilon=24/255$ 9.445 M $\pm$ 0.040 M 27646 MB $\pm$ 614 MB 53.16 min $\pm$ 0.32 min
-unconstrained 12.830 M $\pm$ 0.059 M 77141 MB $\pm$ 1444 MB 73.04 min $\pm$ 0.28 min
-room clean 1.513 M $\pm$ 0.004 M 12316 MB $\pm$ 13 MB 25.36 min $\pm$ 0.09 min
-$\epsilon=8/255$ 3.034 M $\pm$ 0.003 M 19542 MB $\pm$ 363 MB 32.99 min $\pm$ 0.34 min
-$\epsilon=16/255$ 7.129 M $\pm$ 0.020 M 46238 MB $\pm$ 802 MB 47.98 min $\pm$ 0.72 min
-$\epsilon=24/255$ 9.724 M $\pm$ 0.053 M 80142 MB $\pm$ 249 MB 56.87 min $\pm$ 0.51 min
-unconstrained 16.019 M $\pm$ 0.025 M 57540 MB $\pm$ 3983 MB 76.25 min $\pm$ 0.14 min
-stump clean 4.671 M $\pm$ 0.070 M 14135 MB $\pm$ 218 MB 27.06 min $\pm$ 0.31 min
-$\epsilon=8/255$ 5.152 M $\pm$ 0.072 M 15342 MB $\pm$ 121 MB 28.21 min $\pm$ 0.31 min
-$\epsilon=16/255$ 10.003 M $\pm$ 0.292 M 25714 MB $\pm$ 1806 MB 38.48 min $\pm$ 0.59 min
-$\epsilon=24/255$ 14.292 M $\pm$ 0.050 M 34246 MB $\pm$ 1108 MB 46.89 min $\pm$ 0.49 min
-unconstrained 13.550 M $\pm$ 1.257 M 36181 MB $\pm$ 3769 MB 51.51 min $\pm$ 1.90 min
-Appendix D Notations
+Table 6: Black-box attack results on Mip-Splatting (Yu et al., 2024) as the victim system, which can further demonstrate the robust generalization ability of our attack against unknown black-box victim systems.
+Constrained Black-box Poison-splat attack with $\epsilon=16/255$ against Mip-Splatting
+Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes]
+Scene Setting clean poisoned clean poisoned clean poisoned
+NS-chair 0.372 M 1.088 M (2.92$\uparrow$) 6318 33042 (5.23$\uparrow$) 7.85 13.75 (1.75$\uparrow$)
+NS-drums 0.429 M 0.773 M (1.80$\uparrow$) 7474 22036 (2.95$\uparrow$) 8.15 12.60 (1.55$\uparrow$)
+NS-ficus 0.234 M 0.420 M (1.79$\uparrow$) 6144 12374 (2.01$\uparrow$) 6.35 8.97 (1.41$\uparrow$)
+NS-hotdog 0.200 M 1.437 M (7.18$\uparrow$) 5350 58878 (11.01$\uparrow$) 7.08 20.08 (2.84$\uparrow$)
+NS-lego 0.307 M 1.127 M (3.67$\uparrow$) 6696 48682 (7.27$\uparrow$) 7.28 16.43 (2.26$\uparrow$)
+NS-materials 0.216 M 0.505 M (2.34$\uparrow$) 5278 9674 (1.83$\uparrow$) 6.72 8.88 (1.32$\uparrow$)
+NS-mic 0.268 M 0.557 M (2.08$\uparrow$) 5338 19758 (3.70$\uparrow$) 9.12 12.23 (1.34$\uparrow$)
+NS-ship 0.330 M 1.793 M (5.43$\uparrow$) 6500 61026 (9.39$\uparrow$) 8.75 21.05 (2.41$\uparrow$)
+MIP-bicycle 8.683 M DoS 80614 DoS 47.57 DoS
+MIP-bonsai 1.670 M 13.016 M (7.79$\uparrow$) 30876 80826 (2.62$\uparrow$) 23.72 61.82 (2.61$\uparrow$)
+MIP-counter 1.493 M 8.329 M (5.58$\uparrow$) 19478 79904 (4.10$\uparrow$) 26.25 56.75 (2.16$\uparrow$)
+MIP-flowers 3.834 M 7.281 M (1.90$\uparrow$) 50922 80286 (1.58$\uparrow$) 32.85 41.80 (1.27$\uparrow$)
+MIP-garden 5.828 M 8.677 M (1.49$\uparrow$) 70446 80440 (1.14$\uparrow$) 43.20 54.43 (1.26$\uparrow$)
+MIP-kitchen 2.182 M 10.734 M (4.92$\uparrow$) 37024 81006 (2.19$\uparrow$) 30.63 66.92 (2.18$\uparrow$)
+MIP-room 2.080 M 12.949 M (6.23$\uparrow$) 31616 81130 (2.57$\uparrow$) 27.13 77.83 (2.87$\uparrow$)
+MIP-stump 5.920 M 13.925 M (2.35$\uparrow$) 65882 80480 (1.22$\uparrow$) 34.33 56.43 (1.64$\uparrow$)
+Unconstrained Black-box Poison-splat attack with $\epsilon=\infty$ against Mip-Splatting
+Metric Number of Gaussians Peak GPU memory [MB] Training time [minutes]
+Scene Setting clean poisoned clean poisoned clean poisoned
+NS-chair 0.372 M 6.106 M (16.41$\uparrow$) 6318 80732 (12.78$\uparrow$) 7.85 57.73 (7.35$\uparrow$)
+NS-drums 0.429 M 6.818 M (15.89$\uparrow$) 7474 81210 (10.87$\uparrow$) 8.15 67.67 (8.30$\uparrow$)
+NS-ficus 0.234 M 4.847 M (20.71$\uparrow$) 6144 80834 (13.16$\uparrow$) 6.35 45.85 (7.22$\uparrow$)
+NS-hotdog 0.200 M 6.876 M (34.38$\uparrow$) 5350 80630 (15.07$\uparrow$) 7.08 59.80 (8.45$\uparrow$)
+NS-lego 0.307 M 6.472 M (21.08$\uparrow$) 6696 80668 (12.05$\uparrow$) 7.28 64.93 (8.92$\uparrow$)
+NS-materials 0.216 M 5.513 M (25.52$\uparrow$) 5278 81112 (15.37$\uparrow$) 6.72 45.08 (6.71$\uparrow$)
+NS-mic 0.268 M 5.433 M (20.27$\uparrow$) 5338 81020 (15.18$\uparrow$) 9.12 52.83 (5.79$\uparrow$)
+NS-ship 0.330 M 6.848 M (20.75$\uparrow$) 6500 81236 (12.50$\uparrow$) 8.75 66.15 (7.56$\uparrow$)
+MIP-bicycle 8.683 M DoS 80614 DoS 47.57 DoS
+MIP-bonsai 1.670 M DoS 30876 DoS 23.72 DoS
+MIP-counter 1.493 M DoS 19478 DoS 26.25 DoS
+MIP-flowers 3.834 M DoS 50922 DoS 32.85 DoS
+MIP-garden 5.828 M DoS 70446 DoS 43.20 DoS
+MIP-kitchen 2.182 M DoS 37024 DoS 30.63 DoS
+MIP-room 2.080 M DoS 31616 DoS 27.13 DoS
+MIP-stump 5.920 M DoS 65882 DoS 34.33 DoS
+Appendix E Notations
 We provide Table 7 to list all the notations appeared in the paper.
@@
 $\lambda$ Loss trade-off parameters
-Appendix E Visualizations
+Appendix F Visualizations
 In this section, we present both the attacker-crafted poisoned images and the resultant reconstructions by the victim model.
@@
 MIP-counter [Uncaptioned image] [Uncaptioned image] 21.33 dB
+Appendix G Ablation Study on Perturbation Range
+We compare the attack effects (in terms of number of Gaussians, GPU memory and training time) and image degradation (in terms of SSIM and PSNR compared with clean input) under different perturbation ranges $\epsilon=8/255$, $\epsilon=16/255$, $\epsilon=24/255$) on NeRF-Synthetic and MIP-NeRF360 datasets, as shown in Table 10 and Table 11.
+Table 10: Attack performance (number of Gaussians, GPU memory consumption and training time) and image quality (SSIM and PSNR compared with clean images) under different levels of perturbation range on NeRF Synthetic dataset.
+Scene Attack setting Number of Gaussians GPU memory Training time SSIM PSNR
+chair clean 0.494 M 3752 MB 9.45 min - -
+$\epsilon=8/255$ 0.670 M 4520 MB 11.24 min 0.42 34.20
+$\epsilon=16/255$ 0.957 M 8003 MB 13.41 min 0.21 27.94
+$\epsilon=24/255$ 1.253 M 16394 MB 15.89 min 0.16 24.35
+unconstrained 4.152 M 51424 MB 40.71 min 0.07 4.63
+drums clean 0.385 M 3456 MB 8.75 min - -
+$\epsilon=8/255$ 0.469 M 4353 MB 10.19 min 0.45 34.05
+$\epsilon=16/255$ 0.695 M 7190 MB 11.94 min 0.24 27.87
+$\epsilon=24/255$ 0.966 M 14668 MB 14.86 min 0.17 24.36
+unconstrained 4.109 M 62328 MB 43.25 min 0.03 4.49
+ficus clean 0.265 M 3194 MB 6.99 min - -
+$\epsilon=8/255$ 0.246 M 3238 MB 8.28 min 0.39 34.45
+$\epsilon=16/255$ 0.273 M 3783 MB 8.94 min 0.18 28.21
+$\epsilon=24/255$ 0.450 M 8993 MB 11.32 min 0.13 24.53
+unconstrained 3.718 M 50450 MB 36.92 min 0.08 4.41
+hotdog clean 0.185 M 3336 MB 8.57 min - -
+$\epsilon=8/255$ 0.610 M 10362 MB 12.96 min 0.46 33.93
+$\epsilon=16/255$ 1.147 M 29747 MB 17.43 min 0.25 27.56
+$\epsilon=24/255$ 1.553 M 39087 MB 21.26 min 0.17 24.11
+unconstrained 4.272 M 47859 MB 38.85 min 0.07 5.75
+lego clean 0.341 M 3532 MB 8.62 min - -
+$\epsilon=8/255$ 0.469 M 4648 MB 10.38 min 0.48 34.08
+$\epsilon=16/255$ 0.805 M 9960 MB 13.02 min 0.28 27.88
+$\epsilon=24/255$ 1.116 M 18643 MB 15.33 min 0.21 24.30
+unconstrained 4.159 M 78852 MB 42.46 min 0.06 4.78
+materials clean 0.169 M 3172 MB 7.33 min - -
+$\epsilon=8/255$ 0.229 M 3581 MB 8.84 min 0.44 34.25
+$\epsilon=16/255$ 0.410 M 4886 MB 9.92 min 0.21 27.97
+$\epsilon=24/255$ 0.589 M 7724 MB 11.51 min 0.14 24.35
+unconstrained 3.380 M 43998 MB 32.40 min 0.08 5.40
+mic clean 0.205 M 3499 MB 8.08 min - -
+$\epsilon=8/255$ 0.251 M 4038 MB 9.29 min 0.36 34.46
+$\epsilon=16/255$ 0.359 M 6437 MB 10.96 min 0.15 28.17
+$\epsilon=24/255$ 0.514 M 13014 MB 12.89 min 0.10 24.48
+unconstrained 3.940 M 61835 MB 39.02 min 0.08 4.50
+ship clean 0.272 M 3692 MB 8.87 min - -
+$\epsilon=8/255$ 0.516 M 5574 MB 11.01 min 0.55 33.42
+$\epsilon=16/255$ 1.071 M 16666 MB 16.12 min 0.33 27.50
+$\epsilon=24/255$ 1.365 M 29828 MB 18.46 min 0.24 24.03
+unconstrained 4.317 M 80956 MB 44.11 min 0.04 5.31
+Table 11: Attack performance (number of Gaussians, GPU memory consumption and training time) and image quality (SSIM and PSNR compared with clean images) under different levels of constraints on MIP-NeRF360 dataset.
+Scene Attack setting Number of Gaussians GPU memory (MB) Training time (min) SSIM PSNR
+bicycle clean 5.793 M 17748 33.42 - -
+$\epsilon=8/255$ 6.116 M 18608 34.73 0.86 30.80
+$\epsilon=16/255$ 10.129 M 27074 44.44 0.67 26.24
+$\epsilon=24/255$ 13.265 M 34870 51.37 0.52 23.12
+unconstrained 25.268 M 63236 81.48 0.02 6.60
+bonsai clean 1.294 M 10216 21.68 - -
+$\epsilon=8/255$ 1.778 M 11495 21.39 0.79 32.45
+$\epsilon=16/255$ 6.150 M 20564 32.51 0.51 26.83
+$\epsilon=24/255$ 9.321 M 27191 39.84 0.35 23.52
+unconstrained 20.127 M 54506 75.18 0.01 6.27
+counter clean 1.195 M 10750 26.46 - -
+$\epsilon=8/255$ 1.739 M 15133 28.06 0.80 32.24
+$\epsilon=16/255$ 4.628 M 29796 37.63 0.52 26.78
+$\epsilon=24/255$ 6.649 M 47607 43.68 0.35 23.45
+unconstrained 11.167 M 80732 62.04 0.01 6.64
+flowers clean 3.508 M 12520 25.47 - -
+$\epsilon=8/255$ 3.457 M 12391 25.53 0.87 29.18
+$\epsilon=16/255$ 5.177 M 16338 30.05 0.71 25.60
+$\epsilon=24/255$ 7.280 M 20264 34.25 0.58 22.79
+unconstrained 18.075 M 45515 62.62 0.02 6.79
+garden clean 5.684 M 17561 33.19 - -
+$\epsilon=8/255$ 5.122 M 16498 32.17 0.84 29.85
+$\epsilon=16/255$ 7.697 M 23115 40.82 0.68 25.87
+$\epsilon=24/255$ 10.356 M 28130 48.88 0.54 22.82
+unconstrained 21.527 M 52140 83.81 0.04 7.43
+kitchen clean 1.799 M 12442 26.61 - -
+$\epsilon=8/255$ 2.675 M 13760 31.39 0.83 31.58
+$\epsilon=16/255$ 6.346 M 20902 43.24 0.60 26.52
+$\epsilon=24/255$ 9.445 M 27646 53.16 0.45 23.18
+unconstrained 12.830 M 77141 73.04 0.03 7.76
+room clean 1.513 M 12316 25.36 - -
+$\epsilon=8/255$ 3.034 M 19542 32.99 0.75 32.41
+$\epsilon=16/255$ 7.129 M 46238 47.98 0.45 26.64
+$\epsilon=24/255$ 9.724 M 80142 56.87 0.29 23.32
+unconstrained 16.019 M 57540 76.25 0.01 6.42
+stump clean 4.671 M 14135 27.06 - -
+$\epsilon=8/255$ 5.152 M 15342 28.21 0.84 31.20
+$\epsilon=16/255$ 10.003 M 25714 38.48 0.65 25.67
+$\epsilon=24/255$ 14.292 M 34246 46.89 0.49 22.98
+unconstrained 13.550 M 36181 51.51 0.02 7.14
+Appendix H Ablation Study on Poison Ratio
+In the main paper, we conducted all other experiments with 100% poison rate, based on the assumption that attacker can fully control data collection process.
+This assumption is reasonable for most 3D tasks, since the datasets typically are the same scene or object captured from different angles.
+To gain deeper insights into the Poison-splat attack, we explore how varying levels of data poisoning ratio can affect the attack’s effectiveness.
+We conduct experiments using poison ratios of 20%, 40%, 60%, and 80%, and report the number of Gaussians, GPU memory consumption, and training time in Table 12.
+In these experiments, the poisoned views are randomly selected.
+To mitigate the effects of randomness, we report the mean value and standard deviation for each setting across three individual runs.
+Table 12: Attack effectiveness under different poison ratio.
+Scene Poison Ratio Number of Gaussians GPU memory (MB) Training time (min)
+NS-hotdog-eps16 clean 0.185 M $\pm$ 0.000 M 3336 MB $\pm$ 11 MB 8.57 min $\pm$ 0.30 min
+20% 0.203 M $\pm$ 0.004 M 3286 MB $\pm$ 70 MB 8.84 min $\pm$ 0.27 min
+40% 0.279 M $\pm$ 0.024 M 3896 MB $\pm$ 286 MB 9.79 min $\pm$ 0.53 min
+60% 0.501 M $\pm$ 0.054 M 7367 MB $\pm$ 1200 MB 11.33 min $\pm$ 0.82 min
+80% 0.806 M $\pm$ 0.018 M 13621 MB $\pm$ 861 MB 14.02 min $\pm$ 0.60 min
+100% 1.147 M $\pm$ 0.003 M 29747 $\pm$ 57 MB 17.43 min $\pm$ 0.61 min
+NS-ship-eps16 clean 0.272 M $\pm$ 0.001 M 3692 MB $\pm$ 52 MB 8.87 min $\pm$ 0.44 min
+20% 0.321 M $\pm$ 0.004 M 3764 MB $\pm$ 70 MB 9.55 min $\pm$ 0.17 min
+40% 0.472 M $\pm$ 0.021 M 4796 MB $\pm$ 293 MB 10.49 min $\pm$ 0.58 min
+60% 0.718 M $\pm$ 0.003 M 7034 MB $\pm$ 407 MB 12.52 min $\pm$ 0.45 min
+80% 0.924 M $\pm$ 0.004 M 11850 MB $\pm$ 510 MB 14.37 min $\pm$ 0.60 min
+100% 1.071 M $\pm$ 0.003 M 16666 MB $\pm$ 548 MB 16.12 min $\pm$ 0.42 min
+MIP-counter-eps16 clean 1.195 M $\pm$ 0.005 M 10750 MB $\pm$ 104 MB 26.46 min $\pm$ 0.57 min
+20% 1.221 M $\pm$ 0.005 M 11043 MB $\pm$ 141 MB 27.41 min $\pm$ 0.85 min
+40% 1.358 M $\pm$ 0.030 M 11535 MB $\pm$ 147 MB 27.97 min $\pm$ 0.47 min
+60% 2.005 M $\pm$ 0.056 M 13167 MB $\pm$ 227 MB 28.87 min $\pm$ 0.52 min
+80% 3.273 M $\pm$ 0.119 M 19578 MB $\pm$ 1417 MB 32.99 min $\pm$ 0.57 min
+100% 4.628 M $\pm$ 0.014 M 29796 MB $\pm$ 187 MB 37.63 min $\pm$ 0.48 min
+MIP-room-eps16 clean 1.513 M $\pm$ 0.004 M 12316 MB $\pm$ 13 MB 25.36 min $\pm$ 0.09 min
+20% 1.520 M $\pm$ 0.015 M 12624 MB $\pm$ 502 MB 26.45 min $\pm$ 0.75 min
+40% 1.938 M $\pm$ 0.036 M 14495 MB $\pm$ 320 MB 27.99 min $\pm$ 0.72 min
+60% 3.589 M $\pm$ 0.096 M 18760 MB $\pm$ 765 MB 34.09 min $\pm$ 0.79 min
+80% 5.503 M $\pm$ 0.017 M 28166 MB $\pm$ 457 MB 41.47 min $\pm$ 1.36 min
+100% 7.129 M $\pm$ 0.020 M 46238 MB $\pm$ 802 MB 47.98 min $\pm$ 0.72 min
+From Table 12, we have following observations:
+  • Attack performance, in terms of both GPU memory usage and training time extension, becomes stronger as poisoning rate increases.
+  • At higher poisoning rates (between 60%-80%), the attack exhibits greater variance due to the randomness involved in selecting which views to poison.
+This suggests that the choice of views to poison can also impact the effectiveness of the attack.
+Appendix I Ablation Study on Defense Threshold of limiting Gaussian numbers
+In the main paper, we set the defense threshold to ensure a tight wrapping of necessary Gaussians for reconstruction on the clean scene.
+More specifically, for three scenes we show in Section 4.4 and Figure 5, we set defense threshold as in the following Table 13:
+Table 13: Number of Gaussians on clean data and as defense threshold used in Section 4.4.
+Scene #Gaussians on clean data #Gaussians as defense threshold
+Bicycle 5.793 Million 6 Million
+Flowers 3.508 Million 4 Million
+Room 1.513 Million 2 Million
+For a 3DGS trainer, setting a uniform defense threshold is a notably challenging task due to the significant variance in different scenes.
+Implementing a low threshold can safeguard the victim’s computational resources from excessive consumption, but it may also substantially compromise the quality of reconstruction for complex scenes.
+Choosing the defense threshold should consider this resource-quality tradeoff.
+Taking room of Nerf-Synthetic as an example, we tested the defense threshold acrosss 2 million to 7 million, as shown in the following Table 14 and Figure 6:
+Table 14: Defense threshold impact on GPU consumption and reconstruction PSNR.
+Defense threshold Max GPU memory (MB) Reconstruction PSNR
+Clean input 12316 MB 29.21
+2 Million 14192 MB 19.03
+3 Million 16472 MB 22.27
+4 Million 18784 MB 24.11
+5 Million 23642 MB 24.49
+6 Million 36214 MB 28.02
+Poison + No Defense 46238 MB 29.08
+Figure 6: Rendering of reconstructed models under different defense threshold.
+As the defense threshold becomes tighter, the maximum GPU memory usage is confined within a safer range.
+However, this results in a significant drop in the reconstruction’s PSNR, falling to as low as 19.03, which nearly makes it unusable.
+(c.f.
+Figure 6 (A)).
+What’s more, while setting a threshold at 6 million Gaussians appears adequate (achieving PSNR 28.02 dB), this same threshold yields poor rendering results for bicycle scene, with only 20.33 dB PSNR (c.f.
+Figure 6 (F)).
+Overall, determining a universal defense threshold that preserves reconstruction quality across various scenes is challenging.
+Designing a more effective defense against computation cost attack remains an open question, and we hope that future research will continue to explore this area.
+Appendix J Image smoothing is not an ideal defense
+We will show that image smoothing as a pre-processing to Gaussian Splatting training procedure is not an ideal defense.
+Without a reliable detection method, defenders may only resort to universally applying image smoothing to all incoming data.
+This pre-processing will significantly compromise reconstruction quality.
+We found that although image smoothing may reduce GPU consumption to some extent, it severly undermines efforts to preserve fine image details (Yu et al., 2024; Ye et al., 2024; Yan et al., 2024b).
+As illustrated in Figure 7, applying common smoothing techniques such as Gaussian filtering or Bilateral filtering (Tomasi & Manduchi, 1998) leads to a substantial degradation in reconstruction quality.
+For instance, on the chair scene of NeRF-Synthetic dataset, reconstruction achieves 36.91 dB PSNR without pre-processing; however, with Gaussian or Bilateral filtering, the PSNR drops sharply to around 25 dB.
+This level of degradation is clearly undesirable.
+Given these challenges, we urge the community to develop more sophiscated detection and defense mechanisms against computation cost attacks, balancing resource consumption with the preservation of image quality.
+Figure 7: Using image smoothing as a pre-processing step can result in unsatisfactory model reconstruction quality, particularly regarding fine details in images.
+This naive defense strategy significantly degrades the PSNR by approximately 10 dB, contradicting the primary goal of the service provider to maintain high-quality outputs.
+Appendix K Time efficiency of Poison-Splat Attack
+According to Algorithm 1, we need to solve a bi-level optimization problem, and the time complexity is related to $T$ (inner iterations) and $\tilde{T}$ (outer iterations).
+However in practice, our inner iterations $T$ are far fewer than outer iterations $\tilde{T}$, and the actual time cost of attacker is even less than victim’s training time, which is an advantage of our attack.
+Specifically, the total running time of algorithm 1 is decided by two parts:
+  1. 1.
+    Train a proxy model on clean data (Line 2 of Algorithm 1);
+  2. 2.
+    Solve the bi-level optimization (Line 3 to Line 13 of Algorithm 1)
+For example, for unconstrained attacks on MIP-NeRF360, we set $T=6000$ and $\tilde{T}=25$, and the time of attack is totally acceptable, and is even shorter than the victim training, as shown in the Table 15 below:
+Table 15: Comparison of training times for the attacker and the victim.
+Scene Attacker: Proxy training + Bi-level optimization Victim: Training time
+Bicycle 33.42 + 16.57 min 81.48 min
+Bonsai 21.68 + 14.00 min 75.18 min
+Counter 26.46 + 15.23 min 62.04 min
+Flowers 25.47 + 15.05 min 62.62 min
+Garden 33.19 + 15.07 min 83.81 min
+Kitchen 26.61 + 14.77 min 73.04 min
+Room 25.36 + 16.50 min 76.25 min
+Stump 27.06 + 15.63 min 51.51 min
+Lastly, the primary motivation of our work is to expose the severity of this security backdoor, and attack efficiency is beyond the scope of this paper.
+We hope to leave the improvement of attack efficiency in future studies.